### v1

# Dissipating quartets of excitations in a superconducting circuit

Abstract
Over the past decade, autonomous stabilization of bosonic qubits has emerged as a promising approach for hardware-efficient protection of quantum information. However, applying these techniques to more complex encodings than the Schrödinger cat code requires exquisite control of high-order wave mixing processes. The challenge is to enable specific multiphotonic dissipation channels while avoiding unintended non-linear interactions. In this work, we leverage a genuine six-wave mixing process enabled by a near Kerr-free Josephson element to enforce dissipation of quartets of excitations in a high-impedance superconducting resonator. Owing to residual non-linearities stemming from stray inductances in our circuit, this dissipation channel is only effective when the resonator holds a specific number of photons. Applying it to the fourth excited state of the resonator, we show an order of magnitude enhancement of the state decay rate while only marginally impacting the relaxation and coherence of lower energy states. Given that stray inductances could be strongly reduced through simple modifications in circuit design and that our methods can be adapted to activate even higher-order dissipation channels, these results pave the way toward the dynamical stabilization of four-component Schrödinger cat qubits and even more complex bosonic qubits.
preprint: APS/123-QED
I Introduction
Quantum Error Correction (QEC) relies on the measurement of error syndromes to detect spurious evolutions of a physical system embedding quantum information [1]. For stabilizer codes [2], the regular approach consists in measuring a set of stabilizers and decoding the measurement record to apply corrective feedback or feedforward controls. Alternatively, QEC can be made autonomous by engineering adequate dissipation channels [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Autonomous QEC is only practical in the most simple decoding cases, in particular when each feedback control depends on the instantaneous value of a single stabilizer only. This is the case for most single-mode bosonic codes [18], in which a qubit encoded in an oscillator is stabilized by a few operators only. This approach was pioneered in the setting of two-legged cat qubits [8, 19, 20, 21, 22, 23, 24, 25]. The two-legged cat code manifold is spanned by two coherent states |±$|\pm\alpha\rangle$ of a target oscillator $a$. It is stabilized by a Lindblad dynamics entailed by a single dissipation channel $\dot{\bm{\rho}}_{a}=\gamma_{2}\mathcal{D}[{\bf a}^{2}-\alpha^{2}]\bm{\rho}_{a}$, where $\gamma_{2}$ is the dissipation rate and $\mathcal{D}[{\bf L}]\bm{\rho}={\bf L}\bm{\rho}{\bf L}^{\dagger}-\big{(}{\bf L}^{\dagger}{\bf L}\bm{\rho}+\bm{\rho}{\bf L}^{\dagger}{\bf L}\big{)}/2$. The Lindblad operator ${\bf a}^{2}-\alpha^{2}$ is of order two in the annihilation and creation operators $({\bf a},{\bf a^{\dagger}})$ and by extension, we characterize the stabilizing dissipation channel as order-two. It may be engineered by activating a two-to-one photon exchange interaction between the target oscillator and an auxiliary mode $b$, combined with a resonant drive on $b$ to yield the Hamiltonian ${\bf H}_{2}=g_{2}({\bf a}^{2}-\alpha^{2}){\bf b^{\dagger}}+\mathrm{h.c.}$ If the auxiliary mode dissipates excitations into a cold environment at rate $\kappa_{b}$ much larger than the two-to-one photon exchange rate $g_{2}$, the multimode dynamics reduces to the desired single-mode Lindblad dynamics of the target mode with $\gamma_{2}=4g_{2}^{2}/\kappa_{b}$ [26, 27] [^1] This hierarchy is needed to adiabatically eliminate the mode $b$ from the dynamics, but does not need to be respected for the cat manifold to be stabilized [27]. Unless the two mode resonance frequencies are precisely targeted to reach $\omega_{b}=2\omega_{a}$ [24], the interaction needs to be parametrically modulated at $2\omega_{a}-\omega_{b}$ to become resonant. This can be realized with a variety of Josephson circuits such as the one pictured in Fig. 1a, all relying on the four-wave mixing capabilities of Josephson junctions. This strategy was successfully deployed to robustly suppress bit-flips of two-legged cat qubits. On the down side, this code does not offer any protection against phase-flips entailed by single-photon loss in the target mode, which is the dominant noise channel in most experimental platforms.
In order to generalize the approach to codes whose states have more complex structures in phase-space and offer some measure of protection against phase-flips, one typically needs higher order dissipation channels and thus higher order interactions. In particular, the order-four channel corresponding to the Lindblad dissipator $\gamma_{4}\mathcal{D}[{\bf a}^{4}-\alpha^{4}]$ stabilizes a manifold of four coherent states, from which one can single out a sub-manifold of given parity to define a four-legged cat qubit [8]. This qubit is protected against bit-flips as its two-legged counterpart, and single-photon loss events can be detected without perturbing the encoded information by monitoring the photon number parity [8, 29, 30, 31, 32], thus suppressing phase-flips at first-order. Note that even higher order dissipation channels can correct parity flips autonomously [33] or stabilize other rotation-symmetric codes [34] and translation-symmetric codes [35, 14, 16], offering a more robust protection of the encoded information. In earlier experimental works, high-order dissipation channels were obtained from low-order interactions. Indeed, autonomous parity recovery was achieved by combining dispersive interactions and a set of single and two-photon drives [33], autonomous stabilization of the GKP code was achieved by combining Rabi interactions and sequences of single-photon drives [14, 16, 36], and a four-to-two photon exchange interaction—which could be used to engineer a four-photon dissipation channel—was obtained by cascading off-resonant two-to-one photon exchange interactions in a Raman configuration [37] [^2] In that work, the auxiliary mode was not designed to dissipate excitations at a sufficiently large rate to turn the interaction into an effective four-photon dissipation channel.. Unfortunately, this approach can lead to the opening of spurious dissipation channels entailed by each low-order primitive involved in the generation of effective high-order interactions. In principle, the problem can be addressed using a biased-noise auxiliary mode [39, 40, 41], error-transparent interactions [42, 43, 44, 45, 25] or carefully controlling the noise spectrum of the dissipative environment [46], but the experimental implementation of these ideas is a major challenge.
Figure 1: Dissipation engineering in a Josephson circuit. a). A target oscillator (coined the memory, blue) is coupled to an auxiliary dissipative mode (coined the buffer, orange) by a SQUID participating in both modes. When the circuit loops are biased with magnetic fluxes $\Phi_{J}$ and $\Phi_{L}$ following Eq. (1) with $\epsilon(t)$ a pump oscillating at $4\omega_{a}-\omega_{b}$, the SQUID mediates a Kerr-free four-to-one photon exchange interaction, opening a four-photon dissipation channel acting on the memory. The mode impedances $Z_{a,b}$ set the strength of the interaction and are of the order of $R_{q}/10=650~{}\Omega$ in our experiment. b) Energy level-diagram featuring the four-to-one photon pump (purple wriggled arrow) and the buffer single-photon dissipation (orange wriggled arrow). In presence of strong non-linearity, energy levels are unevenly spaced and the pump drives a single transition only (dashed purple arrow represents an off-resonantly driven transition). c) Schematic representation of the circuit used in our experiment. A three-pad superconducting structure hosts the memory and buffer modes with respective anti-symmetric and symmetric voltages across the pads. A distributed $\lambda/2$ resonator capacitively coupled to the structure forms a third mode employed for readout. Foster synthesis of the linear circuit connected to the SQUID yields an effective circuit as in (a) but for the presence of a third mode and for spurious inductances in the arms of the SQUID, responsible for residual non-linearities (see Fig. 2a). Magnetic biases are applied through two flux lines (top). A charge line (bottom) is capacitively connected to the circuit and enables linear drives and single-photon dissipation of the buffer. All lines are filtered by periodic loading to limit single-photon dissipation in the memory [47]. d) Optical microscope image of the sample, featuring all the structures pictured in (c) with matching false colors, surrounded by a ground plane (brown).
An alternative strategy circumventing these pitfalls consists in directly engineering high-order dissipation channels from high-order interactions. In this work, we take this approach and aim at engineering a four-photon loss channel—a key primitive to stabilize four-legged cat qubits—through genuine high-order interactions. Adapting the methods developed to stabilize two-legged cat qubits [21], we proceed by activating a resonant four-to-one photon exchange interaction between a long-lived superconducting resonator (coined the memory) and an auxiliary dissipative mode (coined the buffer). In the idealized setting depicted in Fig. 1a, two $LC$ oscillators are connected in series with a Superconducting Quantum Interference Device (SQUID). The SQUID enables mixing of four waves from the memory $a$, one wave from the buffer $b$, and one pump wave providing the energy needed for the interaction to be resonant. In detail, the circuit features two superconducting loops which are biased with magnetic fields
$\Phi_{J}(t)=\Phi_{0}\big{(}\frac{1}{2}-\frac{\epsilon(t)}{\pi}\big{)}\qquad\qquad\Phi_{L}(t)=\Phi_{0}\frac{\epsilon(t)}{2\pi}$ (1)
where $\Phi_{0}=h/(2e)$ is the flux quantum. The AC signal $\epsilon$, referred to as the pump, is much smaller than 1. In this regime, we express the Josephson potential of the SQUID as a function of the reduced phase $\varphi$ across the modes as $V_{\mathrm{SQUID}}(\varphi)=-E_{J}\cos(\varphi+2\pi\frac{\Phi_{L}}{\Phi_{0}})-E_{J}\cos(\varphi+2\pi\frac{\Phi_{L}+\Phi_{J}}{\Phi_{0}})\simeq 2E_{J}\epsilon(t)\sin(\varphi)$. We note that the potential is $\Phi_{0}$-periodic in $\Phi_{J}$ and $\Phi_{L}$, and offsetting $\Phi_{L}$ by $\Phi_{0}/2$ flips its sign. Expressed in terms of the modes creation and annihilation operators, the phase operator reads $\bm{\varphi}=\varphi^{ZPF}_{a}({\bf a}+{\bf a}^{\dagger})+\varphi^{ZPF}_{b}({\bf b}+{\bf b}^{\dagger})$ where the vacuum fluctuations $\varphi^{ZPF}_{i}$ of each mode are set by the mode impedance following $\varphi^{ZPF}_{i}=(\pi\frac{Z_{i}}{R_{q}})^{\frac{1}{2}}$ and $R_{q}\simeq 6.5~{}\mathrm{k}{\Omega}$ is the resistance quantum [48, 49]. We can then write the circuit Hamiltonian
${\bf H}(t)=\hbar\omega_{a}{\bf a}^{\dagger}{\bf a}+\hbar\omega_{b}{\bf b}^{\dagger}{\bf b}+2E_{J}\epsilon(t)\sin(\bm{\varphi})$ (2)
When $\epsilon=0$, the Hamiltonian (2) represents two uncoupled harmonic oscillators. By letting $\epsilon(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$, the SQUID enables a resonant four-to-one photon exchange interaction between the memory and the buffer—stemming from the fifth order term in the sine expansion. In the rotating frame of each mode and applying the Rotating Wave Approximation (RWA), we get [47]
${\bf H}_{4}=\hbar g_{4}:P_{a}({\bf a^{\dagger}}{\bf a})P_{b}({\bf b^{\dagger}}{\bf b})~{}{\bf a}^{4}{\bf b^{\dagger}}:~{}+~{}\mathrm{h.c.}$ (3)
In this expression, :_::\_:: _ : denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ depends strongly on the phase fluctuations of each mode across the SQUID, and $P_{a}$ and $P_{b}$ are polynomials boiling down to unit factors if the SQUID sine potential is truncated to leading order [47]. Crucially, no other non-linear term in the expansion of ${\bf H}(t)$ survives the RWA when the mode frequencies are not commensurable. Therefore, the memory and buffer remain perfectly linear. In presence of strong single-photon dissipation at rate $\kappa_{b}\gg g_{4}$ in the buffer, the system dynamics reduces to that entailed by a four-photon dissipation channel $\frac{4g_{4}^{2}}{\kappa_{b}}\mathcal{D}[{\bf a}^{4}]$ on the memory [26].
When compared to the analogous case where the same circuit is employed to mediate a two-to-one photon exchange interaction [21, 22, 23, 25], several complications arise. First, the process we target involves the mixing of five waves from the memory and buffer, so that the vacuum fluctuations $\varphi^{ZPF}_{a}$, $\varphi^{ZPF}_{b}$ should not be much smaller than 1 for the interaction rate not to be negligible  [^3] We note that approximating the polynomials in Eq. (3) with unit factors is only valid when $|\varphi^{ZPF}_{a}({\bf a}+{\bf a^{\dagger}})|,|\varphi^{ZPF}_{b}({\bf b}+{\bf b^{\dagger}})|\ll 1$ so that, when the four-to-one interaction is employed to autonomously stabilize four-legged cat states with mean photon number $|\alpha|^{2}$, there exists a maximum value of $\varphi^{ZPF}_{a}\alpha$ beyond which the approximation is no longer valid. This can lead to distortion and eventually destabilization of the cat manifold.. Second, in this regime, any deviation from the idealized circuit considered so far may lead to significant spurious non-linearity of the memory. Indeed, irrespective of the exact mechanism breaking the circuit ideality, the leading non-linear term on the memory scales as $\varphi^{{ZPF}^{4}}_{a}$ (Kerr non-linearity) [47]. A consequence of strong non-linearity is that the memory energy levels are no longer evenly spaced, and the frequency of the four-to-one photon transitions $|n+4\rangle_{a}|0\rangle_{b}\leftrightarrow|n\rangle_{a}|1\rangle_{b}$ that one needs to address depends on the memory level index $n$ (see Fig. 1b). The dominant source of non-linearity in our system, which we investigate in Sec. II.3, involves a Cooper-pair pairing mechanism stemming from stray inductance of the SQUID tracks (purple inductor in Fig. 1c). Unfortunately, it is strong enough to prevent the pump from driving more that one specific four-to-one photon transition. Nevertheless, driving the transition $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$, we demonstrate that quartets of excitations can be dissipated at a rate an order of magnitude higher than the single-photon loss rate in our device, as detailed in the remaining of the paper. Given that stray inductances could be significantly reduced by simple design changes in a future iteration of the experiment, our results paves the way for the stabilization of a four-legged cat manifold by dissipation engineering.
II Circuit design and characterization
II.1 Design
The 2D superconducting circuit we design to implement the idealized circuit analyzed in the previous section is pictured in Fig. 1c-d. At its core lie three superconducting islands (blue and orange squares). The top and central islands are bridged by an Asymmetrically Threaded SQUID (ATS) [21] whose shunt superinductor is made of a chain of 25 Josephson junctions for a total nominal inductance of $L=28.5$ nH. A second identical inductor bridges the central and bottom islands. This symmetric structure supports the memory and buffer, with respective anti-symmetric and symmetric voltages across the pads. The SQUID participates strongly in both modes, resulting in strong phase-fluctuations as prescribed in the previous section. A distributed $\lambda/2$ resonator, colored in green and coined the readout mode, is capacitively coupled to the three-island structure with only a weak participation of the ATS. This mode was originally intended to provide a photon number resolving measurement of the memory leveraging the wave-mixing capabilities of the ATS [23]. In practice, the memory and readout modes were found to be dispersively coupled due to circuit non-idealities (see Sec. II.3) such that dispersive measurement was sufficient to characterize the state of the memory (see Sec. III). Performing Foster synthesis of the linear circuit connected to the SQUID [51, 48, 49], one derives the effective circuit pictured in Fig. 2a with the hierarchy $Z_{r}\ll Z_{a}\sim Z_{b}\sim R_{q}/10$. This effective circuit is analogous to the ideal circuit pictured in Fig. 1a but for the presence of a third mode and for the series inductances in the arms of the SQUID. Correspondence between fluxes applied through the ATS loops $(\Phi_{1},\Phi_{2})$ and fluxes biasing the ideal circuit is made through $\Phi_{L}=-\Phi_{1}$ and $\Phi_{J}=\Phi_{1}+\Phi_{2}$. The fluxes are induced by currents threaded through two dedicated feedlines, denoted as flux lines. We here note that AC fluxes are not independently calibrated in our experiment, such that the condition (1) on the ratio of AC fluxes is not necessarily satisfied. As a result, the modes frequency and non-linearity are slightly renormalized by the pump [47]. Finally, a third feedline, denoted as the charge line, is capacitively coupled to the three-pad structure and to the readout resonator. It allows direct linear drive of the modes, and sets the single-photon decay rate of the buffer and readout modes (see Table 1). All feedlines are filtered through sinusoidal modulation of their geometry to limit radiative decay of the memory [47].
Figure 2: Circuit characterization. a) Effective three-mode circuit obtained by Foster synthesis of the linear environment connected to the SQUID and used to reproduce data presented in (b-d). Fitted parameters are given in Table 1. b) Measured and simulated resonance frequency of the readout mode as a function of applied DC magnetic biases. Black pixels indicate that the resonance frequency could not be extracted, either because it lied out of the probed frequency range (encoded in color, simulation results clipped to the same range) or because the fit was unreliable. Two saddle points, labeled $\mathrm{SP}_{1}$ and $\mathrm{SP}_{2}$, are identified by intersecting double-headed arrows. c) Frequency of the memory and buffer low-energy transitions, along cuts of the frequency map intersecting $\mathrm{SP}_{1}$ and $\mathrm{SP}_{2}$ (respectively in bright and dim colors). Cut angles, indicated by double-headed arrows with matching colors in (b), either maximize or minimize curvature. All frequencies are near-identical around both saddle points, indicating that the ATS is near-perfectly balanced. d) First seven transition frequencies of the memory measured at $\mathrm{SP}_{1}$ (empty blue circles). The anharmonicity is positive and decreases with transition index. This spectrum is well captured by our model accounting for stray inductance in series with the junctions of the ATS (filled black circles).
II.2 Modes dispersion against bias fluxes
In order to characterize the system, we record the modes resonance frequencies against DC magnetic fluxes biasing the ATS loops in absence of pump ($\epsilon(t)=0$). We first monitor the readout mode which is directly probed in single-tone spectroscopy through the charge line. Due to the weak participation of the ATS, the mode frequency dispersion and non-linearity are weak enough that it remains visible at most flux biases, yielding the frequency map represented in Fig. 2b. We note that in the regime $E_{J}{\varphi^{ZPF}_{a,b}}^{2}\simeq\hbar\omega_{a,b}$ we operate in (see Table 1), this map qualitatively differs from the egg-box pattern one would naively expect if the SQUID contribution to the system Hamiltonian (2) were a small perturbation. Nevertheless, the perturbative picture is valid nearby the ATS operating points where the SQUID inductive energy vanishes (see Eq. (1) and following comment). We identify three of these points as saddle points of the frequency map, allowing us to relate the direct currents applied through the flux lines to the actual fluxes threading the ATS (Fig. 2b uses calibrated flux coordinates  [^4] The fluxes $\Phi_{1}$ and $\Phi_{2}$ are known up to an integer number of flux quanta and depend on applied currents through an affine transformation. The map is centered at non-zero applied currents, indicating significant flux offsets., and two non-equivalent saddle points are identified by crossing arrows). Next, we record the memory and buffer resonance frequencies—respectively in two-tone spectroscopy leveraging the memory-readout dispersive coupling and in single-tone spectroscopy—around the saddle points. Fig. 2c features each mode resonance frequency (empty circles) when sweeping applied fluxes through the saddle points of the frequency map, along directions of maximum and minimum curvatures (indicated by double-headed arrows with matching colors in Fig. 2b). Strikingly, the modes have near identical resonance frequencies at both saddle points (markers in bright colors corresponding to the saddle point $\mathrm{SP}_{1}$ are aligned with markers in dim colors corresponding to the saddle point $\mathrm{SP}_{2}$), indicating that the Josephson junctions forming the SQUID have a relative energy difference smaller than $0.01~{}\%$ [47]. We attribute this exceptionally low circuit asymmetry to chance in fabrication and junction aging behavior. Indeed, the asymmetry was found to be two orders of magnitude higher in similar circuits that we have characterized, and an order of magnitude higher during a previous cooldown of the same circuit.
II.3 Stray inductances and memory non-linearity
Unexpectedly, when driving the memory on resonance and subsequently probing the readout resonator, we detect a reflected signal that oscillates with the duration and amplitude of the memory drive pulse [47]. Measuring such Rabi oscillations unambiguously reveals a significant non-linearity of the memory, whose transitions can be addressed individually. Driving the memory with sequences of charge drives (to excite single-photon transitions $|n\rangle_{a}\leftrightarrow|n+1\rangle_{a}$) and/or flux drives (to excite three-photon transitions $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ leveraging the wave-mixing properties of the ATS), we reconstruct the memory mode energy level structure at the saddle point (see Fig. 2d). The anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$ is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$. This pattern is not the one expected for non-linearity induced by the small asymmetry of the SQUID (anharmonicity predicted to be orders of magnitude weaker and with opposite sign at both saddle points [47]) and from the non-linearity of the circuit superinductors (negative anharmonicity near constant with $n$ [47]). On the other hand, the pattern is well captured by a simple model accounting for a small stray inductance in series with the junction on each arm of the SQUID (inductive energy $\epsilon_{L}\gg E_{J}$, see Fig. 2a). Indeed, series inductances renormalize the SQUID potential as $\tilde{V}_{\mathrm{SQUID}}(\varphi)=-E_{J}f(\varphi-2\pi\frac{\Phi_{1}}{\Phi_{0}})-E_{J}f(\varphi+2\pi\frac{\Phi_{2}}{\Phi_{0}})$ where $f(\varphi)=\cos(\varphi)-\frac{E_{J}}{4\epsilon_{L}}\cos(2\varphi)+\mathcal{O}((\frac{E_{J}}{\epsilon_{L}})^{2})$ [53, 54]. The $\cos(2\varphi)$ contribution corresponds to the coherent tunneling of pairs of Cooper pairs. It is maximal around the ATS saddle points and yields a positive anharmonicity that decreases rapidly with the level index $n$ due to the effective doubling of the modes phase-fluctuations [53, 47].
Overall, we fit all data presented in Fig. 2b-d with the energy level structure of the effective circuit pictured in Fig. 2a. The fitted parameters are close to design values (see Table 1). In particular, we note that the memory energy spectrum is well reproduced with a $\cos(2\varphi)$ contribution to the SQUID potential at the level of 0.2 % of the standard $\cos(\varphi)$ term. This is an order of magnitude smaller than the values reported in a recent analysis of transmon energy level structure [55] and can be entirely attributed to the mechanism described above involving series inductances of $\Phi_{0}^{2}/(4\pi^{2}\epsilon_{L})=85$ pH, compatible with the estimated geometric inductance of the circuit. Thus, we do not find signatures of intrinsic distortion of the junctions potential as in [55] which may be explained by a thicker oxide barrier in our junctions [47]. Geometric inductances could be significantly reduced in a future iteration of the experiment by placing the ATS shunt inductor outside of the SQUID loop and scaling down the latter from its current $\sim 50\times 50~{}\mathrm{\mu m}^{2}$ dimensions.
Parameter Value Parameter Value
$\varphi_{a}^{ZPF}$ 0.405 $\omega_{a}$ $2\pi\times 4.13$ GHz
$\varphi_{b}^{ZPF}$ 0.312 $\omega_{b}$ $2\pi\times 6.94$ GHz
$\varphi_{r}^{ZPF}$ 0.044 $\omega_{r}$ $2\pi\times 7.45$ GHz
$E_{J}$ $h\times$17.7 GHz $\delta E_{J}$ $h\times$ 0.35 MHz
$\kappa_{b}$ $2\pi\times$2.05 MHz $\kappa_{r}$ $2\pi\times$1.6 MHz
$\kappa_{\phi}$ $2\pi\times$32 kHz $\epsilon_{L}$ $h\times 1920$ GHz
Table 1: Fitted circuit parameters. $\delta E_{J}$ is the difference in Josephson energy between the two junctions of the SQUID, while $E_{J}$ labels the mean energy. Details of the fit procedure are presented as Supplemental materials.
III Activation and characterization of a high-order dissipation channel
As detailed above, an adverse consequence of the strong non-linearity of the memory is that the memory dissipator $\mathcal{D}[{\bf a}^{4}]$ cannot be activated with a single pump microwave. Combining pumps each addressing a specific transition $|n+4\rangle_{a}|0\rangle_{b}\leftrightarrow|n\rangle_{a}|1\rangle_{b}$ could in principle circumvent the problem. It is however technically challenging as the amplitude and phase of each frequency component needs to be precisely adjusted, and was not attempted. We instead address a specific transition, namely $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ and characterize the engineered dissipator $\mathcal{D}[|0\rangle_{a}\langle 4|_{a}]$ so obtained.
III.1 Preparing a Fock state and measuring its occupation
In order to characterize the engineered dissipation channel, we need to prepare the memory in each Fock state $|n\rangle_{a}$  [^5] The spurious non-linear term in the circuit Hamiltonian is a small perturbation so that eigenstates have an overlap close to 1 with unperturbed Fock states. and to monitor the time decay of its population. Both the preparation and measurement are enabled by the same non-linearity preventing us from addressing more than a single four-to-one photon transition. In principle, the former only requires to apply sequences of $\pi$-pulses on the $|k\rangle_{a}\leftrightarrow|k+1\rangle_{a}$ transitions for $k<n$, while the latter is enabled by the dispersive shift of the readout mode resonance frequency conditioned on $n$. However, preparation and measurement of highly excited states are challenging due to the memory anharmonicity and the readout mode dispersive shift both tailing off with the level index $n$. We therefore limit ourselves to preparation and measurement of Fock states up to $n=4$.
In detail, the states are prepared with pulse sequences combining $\pi$-pulses applied on the single-photon transitions $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|2\rangle_{a}$ (direct charge drive) and/or on the three-photon transitions $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|4\rangle_{a}$ (flux drive of the ATS). The latter are favored over sequences of pulses on single-photon transitions as the detuning between neighboring three-photon transitions is larger, allowing for stronger drives. Yet, all pulses have a duration on par with the coherence time of the transitions, so that preparation is imperfect. Nevertheless, only energy states $|k\rangle_{a}$ with $k\leqslant n$ are occupied when preparing $|n\rangle_{a}$. This allows us to subsequently measure the occupation of $|n\rangle_{a}$ up to an unknown proportionality factor, as we now detail.
The principle of this measurement is depicted in Fig. 3a. While the memory mode vacuum state $|0\rangle_{a}$ exerts a dispersive pull onto the readout mode which is significantly different from the pull exerted by excited states, the differential pull exerted by one excited state and the next tails off rapidly. This makes excited states hard to distinguish directly. We circumvent this shortcoming by applying or not a $\pi$-pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition before probing the readout mode on resonance with a long pulse. Since the dynamics of the decay from $|n+3\rangle_{a}$ to the vacuum state is delayed compared to the decay from $|n\rangle_{a}$, we detect a time-shifted signal in the reflected field $s_{n}^{ON}(t)$ detected after the application of the $\pi$-pulse with respect to the field $s_{n}^{OFF}(t)$ detected in absence of the pulse [47]. We integrate the detected fields with a time-envelope $\mathcal{E}(t)$ chosen to maximize the signal-to-noise ratio of the differential signal $S_{n}=\int_{t}\Re\Big{(}\big{(}s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big{)}\mathcal{E}(t)\Big{)}\mathrm{d}t$ between these two situations [47]. Crucially, the $\pi$-pulse only modifies the memory state when $|n\rangle_{a}$ is occupied so that, even though it does not perfectly transfer population from $|n\rangle_{a}$ to $|n+3\rangle_{a}$, the integrated differential signal $S_{n}$ is proportional to the occupation of state $|n\rangle_{a}$ at the instant when the pulse is delivered. Note that we chose to apply the on/off pulse on a three-photon transition rather than a single-photon transition for similar reasons as for state preparation, and to maximize the differential signal.
Before applying these preparation and measurement methods to characterize the memory population dynamics in presence of a pump, we report that some of the memory transitions—driven either during preparation or measurement, involving a single photon or three—were frequently and transiently observed to shift or split [47]. We attribute this effect to hybridization of the memory to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [57, 58]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, the probability that at least one of the memory transitions be affected by the TLS bath was high enough that waiting for the bath to reconfigure in a more favorable configuration (either passively or cycling the system to room temperature) was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the memory transitions away from any resonant coupling with TLS’s [47]. The frequency of this drive is chosen to minimize decoherence induced on the memory, but still leads to degradation of its energy relaxation and dephasing times. The datasets presented in the following section were acquired in presence of such a drive.
III.2 Characterization of the four-photon dissipation channel
Figure 3: Two-mode dynamics under the four-to-one photon pump a) Principle of the memory photon-number resolving measurement. Dispersive pull on the readout mode saturates for high energy states, preventing their direct resolution. To estimate the occupation of $|n\rangle_{a}$, a $\pi$-pulse is applied on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition, modifying the dynamics of its relaxation to the ground state. Subsequently probing the readout mode on resonance, one obtains a signal $s_{n}^{ON}$ which is time-shifted from the signal $s_{n}^{OFF}$ acquired in absence of pulse. Assuming that $|n\rangle_{a}$ is the highest occupied state before the pulse is applied, the integrated differential signal is proportional to its occupation (see text). b) Time-decay of population in state $|4\rangle_{a}$ (measured up to an unknown proportionality factor) as a function of the pump amplitude (encoded in color and referenced by the voltage $V_{p}$ applied to the mixer generating the pump wave). The decay gets faster for increasing pump amplitude, and eventually deviates from an exponential law. c) Buffer reflection spectrum acquired when simultaneously pumping the four-to-one photon transition (spectra offset proportionally to the pump amplitude for readability). The buffer line shifts, broadens and eventually splits for increasing pump amplitudes, revealing a near-strong hybridization of the $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$ states. Datasets (b) and (c) are reproduced by Lindblad master-equation simulations (lines), whose parameters are given in Table 1. For (c), we fit a displacement rate of $\epsilon_{b}=2\pi\times 175$ kHz yielded by the resonant drive applied to the buffer. d) Four-to-one photon exchange rate against pump amplitude, extracted from these simulations. It slightly deviates from the expected linear scaling.
Armed with the preparation and measurement techniques described in the previous section, and after locating the frequency of the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition dressed by the pump itself [47], we investigate the time decay of population prepared in state $|4\rangle_{a}$ under pump (see Fig. 3b). Clearly, the differential signal $S_{4}$—proportional to the occupation of $|4\rangle_{a}$—decays faster as the pump amplitude increases (here referenced by the voltage $V_{p}$ applied to the mixer generating the pump). For the lowest pump amplitudes ($V_{p}\leqslant 0.1~{}$V), the decay is exponential. However, for larger pump amplitudes, we note a deviation from an exponential law, indicating that reducing the joint dynamics of the memory and buffer modes to a single-mode Lindblad dynamics governed by the dissipator $\mathcal{D}[|0\rangle_{a}\langle 4|_{a}]$ is no longer valid. We further investigate this regime with a complementary experiment in which, for the same range of pump amplitudes, we record the buffer reflection spectrum in continuous-wave (see Fig. 3c). As the pump amplitude increases, a dip in amplitude appears in the buffer line, which eventually splits into two almost separated lines. This behavior is expected at the onset of strong hybridization of the $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$ levels, which occurs when $g_{4}\sqrt{4!}\gtrsim\gamma$ with $\gamma$ the intrinsic width of the transition accounting both for homogeneous and inhomogeneous broadening.
These two datasets are quantitatively reproduced (lines in Fig. 3b-c) by Lindblad master equation simulations modeling the memory and buffer interacting via the Hamiltonian (3[47]. The anharmonicity and relaxation rate of each mode are independently calibrated and accounted for in these simulations. We leave as free fit parameters the value of $g_{4}$ for each pump amplitude and the strength of the buffer drive for the whole second dataset (see Supplemental materials for details of the fit procedures and assumptions behind this model). We also adjust in our model a pure dephasing rate of $\kappa_{\phi}=2\pi\times 32$ kHz for the memory, close to the independently measured dephasing rate of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition [47]. Finally, we fit a proportionality factor between the differential signal $S_{4}$ and the occupation of state $|4\rangle_{a}$. The fitted value of $g_{4}$ is reported in Fig. 3d for each pump amplitude. It reaches $2\pi\times 185~{}$kHz for the strongest pump amplitude, corresponding to a normalized amplitude $|\epsilon(t)|=0.035$ with the convention used in the introduction. Note that the scaling of $g_{4}$ slightly deviates from the expected linear dependency against pump amplitude at room temperature, which could be explained by the saturation of the mixer generating the pump pulses.
Figure 4: Four-photon loss channel. Decay rates $\Gamma_{n}$ of the first four excited states of the memory ($1\leq n\leq 4$) in presence of a pump of varying amplitude. $\Gamma_{4}$ is estimated through spectral analysis [47] of the model Lindbladian used to reproduce the datasets presented in Fig. 3b-c. Following an initial quadratic increase with respect to pump amplitude, it saturates near its maximum theoretical value at half the buffer damping rate (red dashed line). $\Gamma_{n}$ for $n\leq 3$ is directly obtained from an exponential fit of the time-decay of population prepared in $|n\rangle_{a}$ (presented as Supplemental materials). Their increase at large pump amplitude is not captured by our model.
In order to quantitatively relate the largest fitted values of $g_{4}$ to four-photon dissipation rates, one needs to go beyond the simple model reduction allowed by adiabatic elimination of the buffer. To this end, we derive from the two-mode Lindblad master equation an evolution matrix encoding the dynamics of the expectation values of a set of operators, among which $|4\rangle_{a}\langle 4|_{a}$ [47]. Through spectral analysis, we then identify the matrix eigenvalue corresponding to the effective decay $\Gamma_{4}$ of the population in state $|4\rangle_{a}$, and report its value in Fig. 4 (filled red circles). Following an initial quadratic increase of $\Gamma_{4}$ against pump amplitude—also predicted by the adiabatic elimination of the buffer—we observe a saturation of $\Gamma_{4}$ near its theoretical maximum value $\kappa_{b}/2$ set by the buffer damping rate [23]. In this regime, engineered four-photon dissipation dominates the dynamics, inducing decay of population from state $|4\rangle_{a}$ at a rate that exceeds that induced by native single-photon dissipation by a factor 10. As a sanity check, we also record, under the same conditions, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ and report the corresponding decay rates $\Gamma_{n}$ in Fig. 4. While these should nominally not be affected by the pump (see Fig. 1b), we observe a small but clear increase with pump amplitude. Such behavior was reported in similar parametrically pumped systems [19, 21], and are intensively investigated [59, 60, 54, 61]. In any case, the spurious increase of $\Gamma_{n}$ for $n<4$ remains an order of magnitude lower than the recorded increase of $\Gamma_{4}$.
IV Conclusion and outlook
We demonstrate the engineering of a four-photon dissipation channel whose rate far exceeds the intrinsic single-photon loss rate of the target oscillator. Our approach relies on a near Kerr-free Josephson mixing element embedded in a high-impedance environment, mediating genuine high-order interactions while inducing only residual non-linearity on the involved modes. Even though this non-linearity is larger than the engineered dissipation rate in our experiment—preventing us from dissipating quartets of photons from more than a specific oscillator state—we show that it dominantly originates from a renormalization of the SQUID inductive potential by stray inductances in its tracks. We thus estimate that simple design changes would decrease the circuit non-linearity down to a level where it can be dynamically canceled [47]. The four-photon dissipation channel would then become state-agnostic, enabling the dissipative stabilization of a four-legged cat manifold.
Looking forward, stabilization of the four-legged cat manifold could be alternated with parity checks [29, 30, 31, 32] or autonomous parity recovery operations [33] to provide a first-order error-corrected qubit. However, tying up with the discussion given in the introduction, these techniques usually rely on low-order interactions to probe the infinite-order parity operator, with the risk of opening spurious dissipation channels corrupting the encoded qubit. Alternatively, one could increase further the memory impedance presented to the SQUID to enable a genuine parity detection compatible with the manifold stabilization [62]. In a similar regime, one could consider stabilizing Schrödinger cat states with more than four components via wave-mixing processes of even higher order than considered in this work, or combining several such processes to stabilize GKP grid states  [63, 64, 65]. These examples illustrate the potential of driven high-impedance circuits for autonomous and robust quantum error-correction.
Acknowledgments
The authors thank S. Jezouin, R. Lescanne, A. Murani, and W.C. Smith for fruitful discussions on circuit design, J. Palomo, A. Pierret, M. Rosticher and J.-L. Smirr for assistance in microfabrication, and M. Mirrahimi, P. Rouchon and L.-A. Sellem for valuable discussions on dissipation engineering. Our circuit was fabricated within the consortium Salle Blanche Paris Centre and at the Collège de France. This work has been funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101042304), by the French ANR-22-PETQ-0003 grant and by the ANR-22-PETQ-0006 grant under the ‘France 2030’ plan.
Data availability
The experimental data and numerical simulations presented in this manuscript are available from the corresponding author upon request.
Author contributions
A. V. designed and fabricated the sample, performed the experiment and analyzed the data under the supervision of P.C.-I. B.B. and L.L. performed complementary calibration measurements and supporting calculations. P.C.-I. and Z.L. conceived the project. M.V., A.D. and P.M. designed the sample holder used in the experiment. All authors contributed in preparing the manuscript.
Correspondence
Any correspondence should be addressed to philippe.campagne-ibarcq@inria.fr.
Supplemental Materials
S1 Circuit design and measurement setup
S1.1 Target parameters and circuit design
Our sample was designed to dynamically stabilize a four-legged cat qubit. The choice of circuit parameters was motivated by this goal. Here, we give a brief summary of the target dynamics stabilizing a four-legged cat and of the target parameter regime.
In an ideal setting, dynamical stabilization of the four-legged cat manifold (spanned by the four coherent states $|+\alpha\rangle,~{}|+i\alpha\rangle,~{}|-\alpha\rangle,~{}|-i\alpha\rangle$) is achieved through the interaction Hamiltonian
${\bf H}_{4}=\hbar g_{4}({\bf a}^{4}-\alpha^{4}){\bf b^{\dagger}}+\mathrm{h.c.}$ (S1)
with a lossy buffer $b$ (single-photon dissipation rate $\kappa_{b}$), here expressed in the rotating frame of both modes [8]. When $\kappa_{b}\gg g_{4}$, this two-mode dynamics reduces to a single-mode Lindblad-type dynamics
$\dot{\bm{\rho}}=\gamma_{4}\mathcal{D}[{\bf a}^{4}-\alpha^{4}]\bm{\rho}$ (S2)
with $\gamma_{4}=4g_{4}^{2}/\kappa_{b}$. The coherent states spanning the cat manifold are attractors of this dynamics. Note that this property remains true when the above hierarchy between $\kappa_{b}$ and $g_{4}$ is not respected [27]. In (S1), the term $\alpha^{4}{\bf b^{\dagger}}+\mathrm{h.c.}$ is trivially implemented with a resonant drive on the buffer, so that we now focus on the engineering of the more challenging four-to-one photon exchange interaction.
In practice, we engineer this interaction using a pumped ATS participating in the memory and buffer (see Fig. 1). After Foster decomposition of the linear circuit connected to the ATS [48, 21], we reach the Hamiltonian (2) that we recall here:
${\bf H}(t)=\hbar\omega_{a}{\bf a^{\dagger}}{\bf a}+\hbar\omega_{b}{\bf b}{\bf b^{\dagger}}+2E_{J}\epsilon(t)\sin(\varphi^{ZPF}_{a}({\bf a}+{\bf a^{\dagger}})+\varphi^{ZPF}_{b}({\bf b}+{\bf b^{\dagger}}))$ (S3)
where we have neglected the weakly participating readout mode and considered an ideal circuit for simplicity (see Sec. S2 for an analysis of non-idealities). Note that this expression is valid in the limit $|\epsilon(t)|\ll 1$ as it results from a linearization of $\sin(\epsilon(t))$ [21].
We now detail the constraints on the parameters of this Hamiltonian and motivate our choice of parameters.
Zero-point fluctuations of modes across the ATS
We first move to the rotating frame of both modes (${\bf a}\rightarrow{\bf a}e^{-i\omega_{a}t}$ and ${\bf b}\rightarrow{\bf b}e^{-i\omega_{b}t}$). Using Baker-Campbell-Hausdorff formula, we can normal order exponential functions following
$\begin{split}e^{i\varphi^{ZPF}_{a}({\bf a}e^{-i\omega_{a}t}+{\bf a^{\dagger}}e^{+i\omega_{a}t})}&=e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}}\sum_{n,m}\frac{(i{\varphi_{a}^{ZPF}})^{n+m}}{n!m!}e^{i(n-m)\omega t}{\bf a^{\dagger}}^{n}{\bf a}^{m}\\ e^{-i{\varphi_{a}^{ZPF}}({\bf a}e^{-i\omega_{a}t}+{\bf a^{\dagger}}e^{+i\omega_{a}t})}&=e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}}\sum_{n,m}\frac{(-i{\varphi_{a}^{ZPF}})^{n+m}}{n!m!}e^{i(n-m)\omega t}{\bf a^{\dagger}}^{n}{\bf a}^{m}\end{split}$ (S4)
with similar formula for $b$.
If the system is pumped at the frequency $\epsilon_{p}(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$ and after selecting resonant terms, we get in the RWA the static Hamiltonian
$\tilde{{\bf H}}_{4}=g_{4}:P_{a}({\bf a^{\dagger}}{\bf a})P_{b}({\bf b^{\dagger}}{\bf b})~{}{\bf a}^{4}{\bf b^{\dagger}}:~{}+~{}\mathrm{h.c.}$ (S5)
where :_::\_:: _ : denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi{\varphi_{a}^{ZPF}}^{4}{\varphi_{b}^{ZPF}}e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}-\frac{{\varphi_{b}^{ZPF}}^{2}}{2}}$ and
$\begin{split}P_{a}(X)=\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{a}^{ZPF}}^{2n}}{n!(n+4)!}X^{n}\\ P_{b}(X)=\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{b}^{ZPF}}^{2n}}{n!(n+1)!}X^{n}\end{split}$ (S6)
Compared to the target Hamiltonian (S1), higher-order terms appear (corresponding to terms indexed by $n\geqslant 1$ in $P_{a}$ and $P_{b}$). These terms renormalize the four-to-one photon exchange rate between states $|k+4\rangle_{a}|l\rangle_{b}$ and $|k\rangle_{a}|l+1\rangle_{b}$ for $(k,l)\neq(0,0)$. Note that in our experiment, the modes are non-linear so that the pump is not resonant with any of these transitions. Therefore, the system dynamics is the same as that yielded by (S1) (see Sec. S2 and Sec. S4.2).
The four-to-one photon exchange rate $g_{4}$ in (S5) depends strongly on ${\varphi_{a}^{ZPF}}$, which should thus be maximized. However, when stabilizing a four-legged cat in a linear system, high-order terms in the sine expansion can have a a significant impact on the dynamics when $\varphi^{ZPF}_{a}\alpha\gtrsim 1$ such that the sine potential cannot be truncated at leading order. In practice, we saw in numerical simulations that beyond this limit, the cat blobs are distorted and eventually become unstable (not shown). In particular, we found that one should limit ${\varphi_{a}^{ZPF}}\leq 0.5$ to stabilize a cat containing up to 10 photons on average. As for high-order terms in $P_{b}$, they should have no impact on the system dynamics when it is close to its steady state [63]. However, the prefactor $e^{-\frac{{\varphi_{b}^{ZPF}}^{2}}{2}}$ limits the four-to-one photon exchange rate which reaches a maximum for $\varphi^{ZPF}_{b}=1$. In practice, designing a circuit with modes whose vacuum fluctuations across the ATS approach 1 is challenging and one gains only marginally beyond $\varphi^{ZPF}_{b}\sim 0.5$, so that we targeted in design ${\varphi_{a}^{ZPF}}=0.45$ and ${\varphi_{b}^{ZPF}}=0.55$.
As for the readout mode which we have neglected so far, we targeted ${\varphi_{r}^{ZPF}}=0.05$ which was deemed sufficient to perform photon-number resolved measurements of the memory leveraging the non-linearity of the ATS  [^6] We considered activating a longitudinal coupling between the memory and buffer by pumping the ATS at $\omega_{b}$, following the strategy presented in Ref. [23]. In practice this measurement scheme proved hard to tune in our system with significant dispersive interaction between modes, and dispersive measurements were sufficient to achieve the same goal while ensuring that the memory anharmonicity remained smaller than its linewidth at all the ATS DC bias points. This last property is convenient to characterize the system by monitoring the readout mode frequency as a function of bias fluxes (see Fig. 2).
Modes resonance frequencies
The mode resonance frequencies are not particularly constrained toward the goal of stabilizing a four-legged cat. However, the resonance frequencies should be sufficiently separated that the feedlines can be filtered to obtain a strong single-photon dissipation on the buffer while retaining a large intrinsic quality factor for the memory. Moreover, we generate the pump microwave with standard microwave equipment (in the $1-20$ GHz range) and should avoid frequency collisions with processes resonant at $n\omega_{a}+m\omega_{b}$ with $(n,m)\neq(4,-1)$. In practice, we targeted $\omega_{a}=2\pi\times 4.5~{}$GHz and $\omega_{b}=2\pi\times 8~{}$GHz. For the resonance frequency of the readout mode, we aimed at the center of the working range of our parametric amplifier (see Fig. S3).
We adjusted the design of the three-pad structure and shunt inductance values (see Fig. 1) to reach these frequencies and phase fluctuations of each mode across the ATS in finite-element electromagnetic simulations (performed with Ansys HFSS and using the method of energy participation ratios [49]).
Josephson junction energy
The four-to-one photon exchange rate $g_{4}$ scales linearly with $E_{J}$ so that naively, one would aim for large $E_{J}$ values to increase the engineered dissipation rate. However, spurious non-linearities stemming from the tunneling of pairs of Cooper pairs through the SQUID scales with $E_{J}^{2}$ (see Sec. S2). Through numerical simulations, we estimated that for the target values of ${\varphi_{a}^{ZPF}}$ and ${\varphi_{b}^{ZPF}}$ given above, the smallest Josephson junctions that we could fabricate while ensuring symmetry of the SQUID at the $\sim 1~{}\%$ level would yield a sufficiently large four-to-one photon exchange rate to enable inflation of a coherent Schrödinger cat state from vacuum [8]). We therefore targeted this minimum value of $E_{J}=h\times$5 GHz.
The target parameters listed above are summarized in Table 1. The value of the buffer single-photon dissipation rate was optimized in numerical simulations to minimize the inflation time of Schrödinger cat states (for $|\alpha|^{2}\sim 4-10$).
Parameter Target value Measured value
${\varphi_{a}^{ZPF}}$ 0.45 0.405
$\varphi_{b}^{ZPF}$ 0.55 0.312
$\varphi_{r}^{ZPF}$ 0.05 0.044
$\omega_{a}$ $2\pi\times 4.5$ GHz $2\pi\times 4.13$ GHz
$\omega_{b}$ $2\pi\times 8$ GHz $2\pi\times 6.94$ GHz
$\omega_{r}$ $2\pi\times 7.5$ GHz $2\pi\times 7.45$ GHz
$E_{J}$ $h\times$5.6 GHz $h\times$17.7 GHz
$\kappa_{b}$ $2\pi\times$10 MHz $2\pi\times$2.05 MHz
Table 1: Targeted and measured system parameters. Note that the estimated Josephson energy in our experiment is three times larger than its target value, entailing strong non-linearity of all modes (see Sec. S2). The achieved dissipation rate for the buffer is also significantly smaller than its nominal values, which limits the engineered four-photon dissipation rate (see Fig. 4).
S1.2 Fabrication process
Wafer preparation: The circuit is fabricated from a 430 $\mu$m-thick wafer of 0001-oriented, double-side epipolished Sapphire C. The sapphire wafer is initially cleaned through a stripping process in a reactive ion etching (RIE) machine, after which it is loaded into a sputtering system. After one night of pumping, we initiate an argon milling cleaning step, followed by the sputtering of 120 nm of niobium. Subsequently, we apply a protective layer of optical resist (S1805), dice the wafer and clean the small chips in solvents.
Circuit patterning: We spin optical resist (S1805) and pattern the large features (ground plane, feedlines and readout resonator) using a laser writer. After development (MF319), we rinse in de-ionized water for 1 min, and etch the sample in SF6 with a 20 s over-etch. Finally, the sample is cleaned for 10 min by sonication in acetone at $50^{\circ}$ C.
Junction patterning: Next, we apply a bilayer of methacrylic acid/methyl methacrylate [MAA EL13] and poly(methyl methacrylate) (PMMA E3), capped with a layer of conductive resist Electra 92 to evacuate charges during lithography. The entire circuit is patterned in a single electron-beam lithography step. The Electra 92 is removed in deionized water. Subsequently, the development takes place in a 3:1 isopropyl alcohol (IPA)/water solution at $6^{\circ}$ C for 90 s, followed by 10 s in IPA.
Aluminum deposition: The chip is then loaded in an e-beam evaporator. We start with a thorough argon ion milling for 2 min at ±$\pm 30^{\circ}$ angles. We then evaporate 35 nm and 70 nm of aluminum, at ±$\pm 30^{\circ}$ angles, separated by an oxidation step in 100 mbar of pure oxygen for 10 min. After evaporating the two layers, the sample is oxidized for 5 min in 100 mbar of oxygen to provide a controlled oxide cap.
Sample mounting: The chip was subsequently glued with PMMA onto a PCB, wire-bonded and mounted into a sample holder. The device was then thermally anchored to the base plate of a Bluefors dilution refrigerator, surrounded by three concentric cans for magnetic and infrared shielding (outer: cryoperm, middle: aluminum, inner: copper, see Fig. S3).
S1.3 Circuit layout details and measurement setup
Figure S1: Optical microscope pictures of the circuit (false colors). The three-pad structure supporting the memory and buffer modes is visible in blue and orange in the zoomed-in picture on the left (corresponding to the purple rectangle in the zoomed-out picture on the right). The readout resonator is colored in green. The width of the center track of the charge line (light blue) and that of the flux lines (mauve) is sinusoidally modulated to prevent radiative decay of the memory (see Fig. S2). The lengths of the flux lines are not exactly matched.
Circuit layout and filtering of the feedlines
An optical microscope image of our circuit is shown in Fig. S1. The core part of the circuit (zoomed-in picture), comprising the three-pad structure hosting the memory and buffer, the ATS and the termination of the feedlines and readout resonator, is made out of evaporated aluminum. The rest of the circuit (zoomed-out picture), comprising the ground plane, the feedlines and the readout resonator is made out of sputtered niobium. Most of the chip is covered by long meandering coplanar waveguide (CPW) transmission lines. The width of the center track of the lines is periodically modulated to limit radiative decay of the memory, which we now detail.
A 1D microwave transmission line is characterized by its inductance $l$ and capacitance $c$ per unit length. If these electrical properties are periodically modulated with a spatial period $2\pi/\eta$, one obtains a microwave photonic crystal featuring a gap in its energy spectrum. More specifically, for a small modulation amplitude $\epsilon\ll 1$ of both the inductance per unit length $l(x)=l(1+\epsilon\cos(\eta x))$ and the capacitance per unit length $c(x)=c(1+\epsilon\cos(\eta x))^{-1}$—yielding a constant wave propagation velocity $v=1/\sqrt{lc}$ and a modulated characteristic impedance $Z(x)=Z(1+\epsilon\cos(\eta x))$—the gap is centered at $\omega_{G}=v\eta/2$ and its width increases with $\epsilon$ [67]  [^7] This formula holds in the more general case where $v$ also varies, albeit after replacing $v$ with its average value.. A consequence of this gapped spectrum is that the line behaves as a stopband filter for incoming microwave signals, whose in-gap isolation increases exponentially with the number of modulation periods. Note that Stepped impedance filters [69], which are the analog of Bragg mirrors used in optics, were previously employed to this same end [70]. They may be seen as photonic crystals with multiple spatial frequencies $(2n+1)\eta$ such that stopbands appear in their transmission at $(2n+1)v\eta/2$. Photonic crystals were also employed in traveling-wave parametric amplifiers to enable phase-matching [71].
Figure S2: Photonic crystal filters. a) Schematic representation of the charge line CPW geometry, featuring the width of its center track and gaps. Metalized areas are colored in gray (the ground planes at the top and bottom extend beyond the graphic frame, and are connected together by wirebonds). The flux lines layout is identical, except for a reduced length (two periods of modulation instead of five). b) Estimated characteristic impedance of the line, which is locally set by the center track width. c) Numerically computed transmission coefficient of the charge line when connected to 50 $\Omega$ ports (blue). The stopband center frequency is set by the line modulation period, and its width by the amplitude of modulation of its impedance. The flux lines, which are shorter, are not as strongly attenuated in-band.
In our sample, all three feedlines are filtered with these photonic crystals. As illustrated in Fig. S2a, the modulation period is $142\pi/\eta=$ 14 mm for a stopband centered at 4.5 GHz (expected memory frequency). The total CPW width is held constant at 200 $\mu$m, and the inner conductor width is sinusoidally varied from $17~{}\mu$m to $195~{}\mu$m. With these parameters, the line impedance is modulated in the range $20-100~{}\Omega$, and is $50~{}\Omega$ after an integer number of periods. The latter property ensures minimal reflections at the far end of the lines (connected to 50 $\Omega$ ports) for out-of-band microwaves. Note that the impedance variation is not small as in the simple picture presented above, and the modulation is not strictly sinusoidal. Owing to the properties of CPW lines, the phase velocity does not depend significantly on the line geometry.
We perform numerical simulations in order to quantitatively estimate the in-band attenuation of the feedlines. In a nutshell, we use a discrete-cell model for the line (cell electrical length much shorter than relevant wavelengths) and compute the transfer matrix for each cell. We then multiply these matrices to obtain the line transfer matrix. We finally translate the result in a scattering matrix assuming that the ports at each end of the line have an impedance of $50~{}\Omega$ [69]. Intuitively, the transmission coefficient so obtained informs us on the extra protection provided by the modulated line as compared to the case where the system is directly connected to a 50 $\Omega$ port. We report the line transmission against probe frequency in Fig. S2c. The charge line, which is more strongly coupled to the circuit modes, is modulated over 5 periods, yielding more than 40 dB of attenuation. For lack of space on the chip, the flux lines are shorter, with only 2 periods of modulation each, yielding about 15 dB of attenuation. We estimate with electromagnetic simulations that these attenuations are sufficient for radiative decay through the lines not to limit the lifetime of the memory. For all lines, the attenuation varies smoothly over the stopband, whose width exceeds 1 GHz. This feature is particularly appealing as filtering is robust against fabrication disorder, which can result in imprecise targeting of the modes frequency (see Table. 1). Note that the buffer, nominally out of band of the filters, is measured to be close to the stopband. This may partly explain the unexpectedly low value of its single-photon dissipation rate (five times smaller than its nominal value).
Figure S3: Wiring diagram Microwave pulses are generated and detected at room temperature by mixing local oscillators (APUAS from Anapico) with waveforms in the $10-200$ MHz range generated by an OPX from Quantum machines (ADC and DAC ports labeled by their function). DC currents biasing the circuit and SNAIL-parametric amplifier [72] (SPA) are provided by voltage sources from Yokogawa and are thermalized with Therma uD RC filters from Aivon anchored on the 4 K plate of the dilution cryostat. Microwave input lines are attenuated and filtered, in particular with infrared filters (IR) from Bluefors. Output signals are amplified by the SPA and HEMT amplifiers from Low-noise factory and other amplifiers at room temperature.
S2 Non-idealities and non-linearity of the modes
Figure S4: Sources of non-linearity. a) Ideal circuit (no source of non-linearity) layout featuring the memory and buffer modes only. b) SQUID imbalance directly appears in the effective circuit found by Foster synthesis of the environment seen by the SQUID. c) Series inductance in the arms of the SQUID enable a two-Cooper pair tunneling mechanism [53] (double-cross junctions) which is not suppressed at the circuit bias point. d) A superinductor made of finite-length chains of Josephson junctions may be viewed as an inductor in parallel of a chain of inductive elements with quartic potential (spider symbols following the convention of Ref. [48]). In general, the effective modes may present a different impedance $\tilde{Z}_{a,b}$ to this element from the imepedance presented to the SQUID. However, given the circuit symmetry, we find $\tilde{Z}_{a,b}=Z_{a,b}$. e) Asymmetric pump waves yield non-linear contributions to the circuit Hamiltonian in the RWA.
Following the simple analysis given in Sec. I, the ATS [21] behaves, when properly biased, as a Kerr-free Josephson mixing element. In this section, we review mechanisms that break this property and estimate the non-linearity induced on the circuit modes.
S2.1 Asymmetry of the SQUID junctions
We consider the effective two-mode circuit pictured in Fig. S4b, which is obtained after Foster decomposition of the environment seen by the SQUID in the physical circuit (Fig. S4a), and neglecting all modes but the memory and buffer. We derive the ATS inductive potential when the junctions forming its SQUID are imbalanced. Following the convention of Fig. S4b, we let $\varphi$ the reduced phase across the modes in series with the ATS, $\Phi_{1}$, $\Phi_{2}$ the fluxes biasing each loop, $E_{J_{1}}$, $E_{J_{2}}$ the energies of each junction. Defining
$\begin{split}\varphi_{\Sigma}&=\frac{\Phi_{1}+\Phi_{2}}{2\varphi_{0}}\\ \varphi_{\Delta}&=\frac{\Phi_{1}-\Phi_{2}}{2\varphi_{0}}\\ E_{\Sigma}&=E_{J_{1}}+E_{J_{2}}\\ E_{\Delta}&=E_{J_{1}}-E_{J_{2}}\end{split}$ (S7)
($\varphi_{0}=\hbar/(2e)$ is the reduced flux quantum) and using Kirchhoff’s law, we find the ATS potential to be
$V^{ATS}_{\varphi_{\Sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\varphi-\varphi_{\Delta})$ (S8)
Now focusing on the ATS operating point at $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,-\pi/2)$ ($\mathrm{SP}_{2}$ in Fig. 2b), the static circuit Hamiltonian (no pump applied here) reads
${\bf H}_{asym}=\sum_{i}\hbar\omega_{i}{\bf a}^{\dagger}_{i}{\bf a}_{i}-E_{\Delta}\cos(\bm{\varphi})$ (S9)
where $i=a,b$ (${\bf a}_{i}$ denoting the annihilation operator of mode $i$), $\bm{\varphi}=\sum_{i}\varphi^{ZPF}_{i}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ and $\varphi^{ZPF}_{i}=\big{(}\pi Z_{i}/R_{q}\big{)}^{1/2}$. The cosine term in this Hamiltonian features non-linear, photon-number preserving terms that survive the RWA. At leading order (truncating the cosine at fourth order), we find that each mode inherits a self-Kerr non-linearity of amplitude $\chi_{aa}=\frac{1}{2}E_{\Delta}\varphi_{a}^{ZPF~{}4}$, $\chi_{bb}=\frac{1}{2}E_{\Delta}\varphi_{b}^{ZPF~{}4}$ and both modes are dispersively coupled through a cross-Kerr term $\chi_{ab}=E_{\Delta}\varphi_{a}^{ZPF~{}2}\varphi_{b}^{ZPF~{}2}$. Given the near perfect symmetry of the ATS in our experiment, these terms are negligible ($\chi_{aa},\chi_{bb},\chi_{ab}<2\pi\times$10 kHz).
S2.2 Stray inductances and multi-Cooper pair tunneling
Figure S5: Inductive dipole made of a junction in series with an inductor. Using Kirchhoff’s law, the dipole potential may be expressed as a function of its generalized phase $\psi$ only.
Following the derivation in Ref. [53], a junction in series with a small inductor (inductive energies $E_{J}\gg\epsilon_{L}$) behaves as an inductive dipole with reduced phase $\psi$ (see Fig. S5). Letting $i$ the current threading the dipole, $\psi_{L}$ and $\psi_{J}$ the reduced phases across the inductor and the junction, we find using Kirchhoff’s law
$\begin{split}\psi_{L}+\psi_{J}=\psi\\ E_{J}\sin(\psi_{J})=\epsilon_{L}\psi_{L}\end{split}$ (S10)
The dipole energy reads $V=-E_{J}\cos(\psi_{J})+\frac{\epsilon_{L}}{2}\psi_{L}^{2}$. Given the hierarchy $E_{J}\gg\epsilon_{L}$, we can solve the equation (S10) to leading order in $E_{J}/\epsilon_{L}$ and inject the solution in the dipole potential to get
$V(\psi)=-E_{J}\cos(\psi)+\frac{E_{J}^{2}}{4\epsilon_{L}}\cos(2\psi)+\mathcal{O}\big{(}\frac{E^{3}_{J}}{\epsilon^{2}_{L}}\big{)}$ (S11)
The first-order correction to the Josephson potential corresponds to the coherent tunneling of pairs of Cooper pairs. Remaining perturbative terms corresponding to tunneling of larger multiplets of Cooper pairs are neglected.
When replacing standard junctions with two such dipoles in the circuit ATS (see Fig. S4c), the inductive potential is renormalized as
$\tilde{V}^{ATS}_{\varphi_{\sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\varphi-2\varphi_{\Delta})$ (S12)
where $e_{\Sigma}=\frac{E^{2}_{J}}{2\epsilon_{L}}$ and we have assumed the SQUID to be balanced for simplicity. Now focusing on the ATS operating point at $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,-\pi/2)$, the static circuit Hamiltonian (no pump applied here) reads
${\bf H}_{2\varphi}=\sum_{i}\hbar\omega_{i}{\bf a}^{\dagger}_{i}{\bf a}_{i}+e_{\Sigma}\cos(2\bm{\varphi})$ (S13)
This term in $\cos(2\bm{\varphi})$ is the dominant contributor to the modes non-linearity at the saddle point. In particular, the memory anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$ is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$ (see Fig. 2b). This last property is a consequence of the large zero point fluctuations of the phase of the mode, which are effectively doubled by the Cooper pair pairing mechanism. Indeed, the $\cos(2\bm{\varphi})$ term cannot be truncated at fourth order, which would have yielded a Kerr non-linearity (constant anharmonicity).
We fit the memory energy spectrum with this model (including corrections from the Josephson junctions asymmetry and a non-linear contribution from the circuit superinductors, see following section), $e_{\Sigma}$ being the only free fit parameter. We thus estimate $e_{\Sigma}$ and finally $\epsilon_{L}=h\times 1920~{}$GHz. This inductive energy corresponds to a series inductance of 85 pH in each arm of the SQUID, which matches a rough estimate of the geometric inductance of the tracks (each arm is 120 $\mu$m long, see Fig. S1). We note that a recent paper [55] reported deviations from the standard $\cos(\varphi)$ potential for junctions embedded in several transmon circuits, with significant higher-order harmonics. This effect could contribute to the $\cos(2\varphi)$ term observed in our experiment. We can nevertheless place a conservative upper bound at 0.2 % for the ratio of the second to first harmonic in the junction potential (corresponding to the case where the $\cos(2\varphi)$ term is entirely attributed to intrinsic distortion of the junction potential). This bound is an order of magnitude smaller than the values reported in [55]. We note that our Josephson junctions are fabricated with a long, high-pressure oxidation step (see Sec. S1.2), probably resulting in an oxide barrier thicker than for the devices analyzed in that work. We can thus conclude, in line with the model presented in [55], that junctions with a thick oxide barrier are less prone the formation of pinholes and are described by a nearly pure $\cos(\varphi)$ potential.
S2.3 Non-linearity induced by the superinductors
The two superinductors in our circuit are made of chains of $N=25$ Josephson junctions, inducing non-linearity in the modes they participate in. In detail, assuming that all junctions of the chain to be identical, that self-resonant modes of the structure are never excited [73] and that phase-slips occur at a negligible rate [74], internal degrees of freedom of the chain are frozen and the phase drop $\psi$ across the chain is equally distributed among the junctions. The chain potential then reads
$V_{\mathrm{chain}}(\psi)=-NE^{0}_{J}\cos(\psi/N)\simeq-\frac{E^{0}_{J}}{2N}\psi^{2}+\frac{E^{0}_{J}}{24N^{3}}\psi^{4}$ (S14)
where $E_{J}^{0}$ is the energy of each junction and we have truncated the potential at fourth order. This approximation is motivated by the large value of $N$ resulting in a small phase drop across each junction.
In order to quantitatively estimate the Kerr non-linearity induced on the modes, we adapt the energy participation ratio method [49]. We perform finite-element electromagnetic simulations to estimate the phase fluctuations $\varphi^{ZPF}_{i,k}$ of each mode $i=a,b,r$ across each superinductor $k=1,2$. Letting $E_{L}=\frac{E^{0}_{J}}{N}$ the inductive energy of each chain, we then estimate that each mode $i$ inherits a self-Kerr non-linearity of amplitude $\chi_{ii}=\sum_{k}\frac{E_{L}}{2N^{2}}{\varphi_{i,k}^{ZPF}}^{4}$ and any pair of modes $(i,i^{\prime})$ are dispersively coupled through a cross-Kerr non-linearity of amplitude $\chi_{ii^{\prime}}=\sum_{k}\frac{E_{L}}{N^{2}}{\varphi_{i,k}^{ZPF}}^{2}{\varphi_{i^{\prime},k}^{ZPF}}^{2}$. We find these terms to have an amplitude in the tens of kHz range, with a negligible impact on the system dynamics. We still include them in numerical simulations.
S2.4 Miscalibration of the ATS flux biases
While the ATS should in principle be driven with symmetric fluxes to activate the desired four-to-one photon exchange (see Sec. I), we actually drive the circuit with an unknown combination of common and differential flux drives (see Sec. S1.2). Dynamical renormalization of the modes non-linearity may then arise.
To understand this phenomenon, we consider the ideal circuit Hamiltonian (S3) (other sources of non-linearity neglected for clarity), at the ATS operating point and in presence of flux drives biasing the loops both symmetrically (reduced amplitude $\epsilon_{c}(t)$) and antisymmetrically (reduced amplitude $-\epsilon_{d}(t)$). In the rotating frame, it reads
${\bf H}_{cd}(t)=E_{\Sigma}\sin\big{(}\epsilon_{c}(t)\big{)}\sin\big{(}\bm{\varphi}(t)+\epsilon_{d}(t)\big{)}$ (S15)
where $\bm{\varphi}(t)=\sum_{j=a,b,r}\varphi^{ZPF}_{j}({\bf a}_{j}e^{-i\omega_{j}t}+{\bf a}_{i}^{\dagger}e^{+i\omega_{j}t})$. Note that we expressed the Hamiltonian in a gauge where the differential drive appears inside the sine term, which may lead to the apparition of an effective charge drive in the Hamiltonian [75]. The amplitude of this term depends on the exact magnetic field distribution with respect to the capacitive structures in the circuit and is hard to predict a priori. Since it does not yield resonant terms in the RWA, we neglect it in our analysis for simplicity. However, we point out that this drive—along with an effective flux drive stemming from the first-order term in the sine expansion—displaces the states of the modes, yielding new dynamical corrections to the modes non-linearity at the second order of the RWA.
Assuming that the drive amplitudes are small, we expand to second order in $\epsilon_{c,d}$ and find
${\bf H}_{cd}(t)=E_{\Sigma}~{}\epsilon_{c}(t)\sin\big{(}\bm{\varphi}(t)\big{)}+E_{\Sigma}~{}\epsilon_{c}(t)\epsilon_{d}(t)\cos\big{(}\bm{\varphi}(t)\big{)}$ (S16)
When driving the four-to-one photon transition, both drives $\epsilon_{c}(t)=\xi_{c}e^{i\omega_{p}t}+\mathrm{c.c}$ and $\epsilon_{d}(t)=\xi_{d}e^{i\omega_{p}t}+\mathrm{c.c}$ are applied at the frequency $\omega_{p}=4\omega_{a}-\omega_{b}$. In the RWA, the first term yields the desired four-to-one photon exchange ${\bf H}_{4}$. As for the second term, it yields an effective static Hamiltonian
${\bf H}_{cd,RWA}=\overline{{\bf H}_{cd}(t)}=2\Re\big{(}\xi_{c}\xi^{\ast}_{d}\big{)}E_{\Sigma}~{}\overline{\cos\big{(}\bm{\varphi}(t)\big{)}}$ (S17)
This effective Hamiltonian is analogous to the non-linear term in the Hamiltonian (S9) deriving from imbalance of the SQUID. In practice, it slightly renormalizes the modes non-linearity, which we do not account for in our simulations (as detailed in Sec. S4.2, the system dynamics under pump does not depend on the exact value of the non-linearity).
We also note that this term is resonant irrespective of the pump frequency. Thus, it should be possible to apply an off-resonant pump ($\omega_{p}\neq 4\omega_{a}-\omega_{b}$) to dynamically cancel the non-linearity stemming from imbalance of the SQUID. One simply adjusts the symmetric and antisymmetric drive amplitudes and phases such that $2\Re\big{(}\xi_{c}\xi^{\ast}_{d}\big{)}=E_{\Delta}$, where $E_{\Delta}$ is the difference of Josephson energy for the two junctions. However, given that the above analysis is only valid in the limit $|\epsilon_{c}|,|\epsilon_{d}|\ll 1$, this method may only prove useful in the limit $E_{\Delta}\ll E_{\Sigma}$. Finally, we note that the effective Hamiltonian (S17) also induces a shift of the modes frequency. This mechanism is expected to contribute to the controlled frequency shift (induced by an off-resonant pump) that we apply to avoid hybridization of our system with two-level systems (see Sec. S5).
S3 Details of the calibration and measurement sequences
S3.1 Photon-number resolving measurement
Figure S6: Reference signals for the photon-number resolving measurement. In order to maximize SNR, we acquire reference signals $r_{n}^{ON}(t)$ and $r_{n}^{OFF}(t)$ after preparing the state $|n\rangle_{a}$ and applying (ON) or not (OFF) a pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition. We use each signal as an averaging envelope for subsequent measurement of the occupation of $|n\rangle_{a}$ for an unknown state (Eq. (S18)). The non-zero signal at long time for $n=0$ is attributed to residual thermal occupation of state $|n=1\rangle_{1}$ at equilibrium. The maximum of each signal is expected to be delayed when $n$ increases, which is clearly visible for $n=4$ (the initial drop is not understood). The pulses duration and amplitude (referenced by the voltage applied to the mixer generating the pulse) are 0.3 $\mu$s / 0.4 V, 1 $\mu$s / 0.15 V, 1 $\mu$s / 0.15 V and 0.5 $\mu$s / 0.1 V for $n=1,~{}2,~{}3$ and $4$ respectively. These characteristics are only optimized to maximize signal for $n=4$, which explains a higher signal than for $n=2,3$. Preparation of $|n=2\rangle_{a}$, which relies on two successive $\pi$-pulses on single-photon transitions is less efficient than the preparation of $|n=3\rangle_{a}$ (single three-photon $\pi$-pulse), explaining a weaker signal.
The principle of this measurement is described in the main text and depicted in Fig. 3a. We recall here the motivation for the measurement sequence and give details on its calibration.
While the memory mode vacuum state $|0\rangle_{a}$ exerts a dispersive pull onto the readout mode significantly different from the pull exerted by excited states, the differential pull exerted by one excited state and the next tails off rapidly. Indeed, the dispersive pulls stem from the $\cos(2\bm{\varphi})$ contribution in the circuit static Hamiltonian (see Sec. S2) and follow the same trend as the memory mode anharmonicity (see Fig. 2). This makes excited states hard to distinguish directly. We circumvent this shortcoming by applying (ON) or not (OFF) a $\pi$-pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition before probing the readout mode on resonance with a long (40 $\mu$s) pulse. The reflected field is demodulated and its quadratures averaged to obtain the complex signals $s_{n}^{ON}(t)$ and $s_{n}^{OFF}(t)$. Since the dynamics of the decay from $|n+3\rangle_{a}$ to the vacuum state is delayed compared to the decay from $|n\rangle_{a}$, $s_{n}^{ON}(t)$ detected after the application of the $\pi$-pulse is distorted and time-shifted with respect to the signal $s_{n}^{OFF}(t)$ detected in absence of the pulse.
In order to maximize the measurement signal-to-noise ratio (SNR), we first prepare $|n\rangle_{a}$ with a sequence of $\pi$-pulses on single or three-photon transitions, and record an averaged reference signal $r_{n}^{ON}(t)$ and $r_{n}^{OFF}(t)$ (see Fig. S6). Subsequently, when measuring an unknown state of the memory, we integrate the detected differential signal with an envelope provided by these reference signals:
$S_{n}=\frac{1}{T^{1/2}_{\mathrm{bin}}}\dfrac{\int_{t=0}^{T}\Re\Big{(}\big{(}s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big{)}\big{(}r_{n}^{ON}(t)-r_{n}^{OFF}(t)\big{)}^{\ast}\Big{)}\mathrm{d}t}{\Big{(}\int_{t=0}^{T}|r_{n}^{ON}(t)-r_{n}^{OFF}(t)|^{2}~{}\mathrm{d}t\Big{)}^{1/2}}$ (S18)
where $T_{\mathrm{bin}}=200$ ns is the integration time for the acquisition of each point of the raw signals $s_{n}^{ON}$, $s_{n}^{OFF}$, $r_{n}^{ON}$, $r_{n}^{OFF}$.
We use this measurement to record the time-decay of various Fock states under pump, and make several key assumptions. First, we assume that before the measurement, only Fock states $|k\rangle_{a}$ with $k\leq n$ are occupied. Moreover, we assume that this property is preserved during the measurement. In other words, we neglect heating mechanisms such as coupling to a hot bath or measurement-induced transitions to higher-excited states. Finally, we assume that the $\pi$-pulse does not impact the system if states $|k\rangle_{a}$ with $k<n$ are occupied (see below). Therefore, even though the $\pi$-pulse duration is of the order or longer than the states lifetimes, it only modifies the memory state when $|n\rangle_{a}$ is occupied beforehand. Moreover, the signal $S_{n}$ is proportional to the occupation probability.
In order to maximize SNR while ensuring that the $\pi$-pulse does not impact the system if only states $|k\rangle_{a}$ with $k<n$ are occupied, we need to adjust finely its amplitude and duration. Intuitively, we want to minimize the pulse duration while ensuring that it is sufficiently weak and well-resolved in frequency not to excite the $|n-1\rangle_{a}\leftrightarrow|n+2\rangle_{a}$ transition. To this end, we prepare in separate experiments the state $|n\rangle_{a}$ and the state $|n-1\rangle_{a}$ and vary the pulse amplitude and duration (square waveform with smoothed edges). We then select the characteristics yielding the highest signal when $|n\rangle_{a}$ is prepared while yielding a negligible signal when $|n-1\rangle_{a}$ is prepared (see Fig. S7). Note that after this calibration, we could again acquire reference signals for the integration of $S_{n}$ using the optimal $\pi$-pulse characteristics. This was done for $S_{4}$, but not for $S_{n}$ with $n\leq 3$ (we have verified that the reference signal shape does not depend significantly on the $\pi$-pulse characteristics as long as it does not excite unwanted transitions).
Figure S7: Optimization of the $\pi$-pulses parameters for the photon-number resolving measurement. The measurement relies on the differential signal obtained when exciting or not the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition. Holding the exciting pulse duration constant, we vary its amplitude and record the differential signal $S_{n}$ after preparing $|n\rangle_{a}$ (plain lines) or $|n-1\rangle_{a}$ (dashed lines). We select the amplitude yielding the strongest signal in the former case while yielding a negligible signal in the latter case (vertical lines). We repeat this calibration for various pulse durations and select the optimal amplitude combination of amplitude and duration.
S3.2 Tuning the pump frequency
Figure S8: Pump frequency calibration. The pump and buffer frequencies dressed by the pump itself are calibrated iteratively (see text). The buffer frequency $\omega_{b_{1}}$ found at step 1 and $\omega_{b_{3}}$ found at step 3 are slightly mismatched at low pump amplitude. This is attributed to inaccuracy of the estimate made at step 3, in which the reflection was not properly fitted but only smoothed aggressively before locating the minimum of the signal magnitude. This mismatch then propagates to the calibrated pump frequency (step 2 vs step 4) as the pump resonance condition reads $\omega_{p_{n}}=4\omega_{a_{n}}-\omega_{b_{n-1}}$ where $\omega_{b_{n-1}}$ is the frequency at which the buffer is driven (found at the previous step) and $\omega_{a_{n}}$ is the memory frequency dressed by the pump at $\omega_{p_{n}}$.
In Fig. 3, we report the decay of population in state $|4\rangle_{a}$ and the buffer reflection spectrum as the amplitude of the pump addressing the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition is varied (encoded in color). For each pump amplitude, the pump frequency needs to be adjusted as the modes resonance frequency gets shifted through by the pump itself (shifting mechanisms include—but are not limited to—miscalibration of the pump symmetry described in Sec. S2 and direct Stark shift owing to the modes non-linearity). This shift depends non trivially on the pump frequency, mostly due to the uncalibrated dispersion of the cryostat input lines. For a given pump amplitude, the pump frequency is adjusted following four steps.
  1. 1.
    We apply the pump at frequency $\omega_{p_{0}}=4\omega_{a}-\omega_{b}$ corresponding to the four-to-one photon transition frequency in absence of pump. We then probe the buffer in reflection and find its resonance frequency $\omega_{b_{1}}$ dressed by the pump.
  2. 2.
    We apply a weak drive at $\omega_{b_{1}}$ to the buffer and sweep the pump frequency until we see the memory excite at $\omega_{p_{2}}$. This occurs when $\omega_{p_{2}}=4\omega_{a_{2}}-\omega_{b_{1}}$, where $\omega_{a_{2}}$ is the memory resonance frequency dressed by the pump at $\omega_{p_{2}}$. Note that $\omega_{b_{1}}$ is slightly detuned from the frequency $\omega_{b_{2}}$ of the buffer dressed by the pump. Photons from the buffer drive are still converted to quartets of excitations in the memory if $|\omega_{b_{1}}-\omega_{b_{2}}|\lesssim\kappa_{b}$.
  3. 3.
    Again as in step (1), we hold the pump at $\omega_{p_{2}}$ and probe the buffer to find its resonance frequency $\omega_{b_{3}}$. Note that the buffer spectra acquired at this step are the ones reported in Fig. 3 and in Fig. S10. For the largest pump amplitudes, the buffer line splits owing to the hybridization of $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$. Fitting to the full two-mode model presented in Sec. S4.2 was not attempted at this step, so that the buffer resonance frequency was simply extracted by smoothing each spectrum over a wide moving window and locating the minimum of the smoothed signal magnitude.
  4. 4.
    Again as in step (2), we apply a weak drive at $\omega_{b_{3}}$ to the buffer and sweep the pump frequency until we see the memory excite at $\omega_{p_{4}}=4\omega_{a_{4}}-\omega_{b_{3}}$, with $\omega_{a_{4}}$ the memory frequency dressed by the pump at $\omega_{p_{4}}$
One should iterate further the procedure until the resonance condition $\omega_{p_{n}}=4\omega_{a_{n}}-\omega_{b_{n}}$ is met self-consistently. In practice, the variation of the pump frequency $|\omega_{p_{2}}-\omega_{p_{4}}|$ between step (2) and (4) was deemed small enough (less than 0.6 MHz at all considered pump amplitudes) that the buffer frequency $\omega_{b_{4}}$ dressed by the the pump at $\omega_{p_{4}}$ would be near identical the frequency $\omega_{b_{3}}$ measured under the pump at $\omega_{p_{2}}$. This assumption, together with the rather coarse estimate of $\omega_{b_{1}}$ in step (3) explains a small detuning between the predictions from our model (plain lines in Fig. S10) to the recorded spectra (dots).
For each pump amplitude, the time-decay of population in state $|4\rangle_{a}$ (presented in Fig. 3b) was recorded after adjusting the pump frequency at $\omega_{p_{2}}$.
S3.3 Decay rate of low energy states under pump
As a sanity check of our experiment, we record, for the same range of pump amplitudes activating the four-photon dissipation, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ (see Fig. S9). For all states, we observe a slight increase of the decay rate induced by the pump, which is not captured by our model. For the states $|2\rangle_{a}$ and $|3\rangle_{a}$, we also record an initial sharp drop of the corresponding signals $S_{2}$ and $S_{3}$. This drop is not understood, but does not seem directly induced by the pump as it is also visible for a null pump amplitude (dark blue). We thus rather suspect an effect caused by the proximity of the memory preparation pulses and of the final pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition (used in the measurement) for the shortest wait times. All signals are fitted with exponential laws and the fitted decay rates are reported in Fig. 4.
Figure S9: Decay rate of the low-energy states under pump. The time decay under pump of population in states $|1\rangle_{a}$ (top), $|2\rangle_{a}$ (middle) and $|3\rangle_{a}$ (bottom) is recorded (filled circles) and fitted with an exponential law (lines). For each state $|n\rangle_{a}$, we fit a single scaling factor for the differential signal $S_{n}$, close to the value calibrated in Fig. S7 (the slightly lower values recorded here, in particular for $S_{3}$, are attributed to drifts during the $\sim 12$ hour acquisition time). The stationary value of the exponential law is fixed to zero, except for $S_{1}$ where we leave it as a free fit parameter to account for thermal excitations of the memory at equilibrium (see Fig. S6). The initial drop of $S_{2}$ and $S_{3}$ is not understood but does not seem directly related to the pump as it is visible at all pump amplitudes.
Figure S10: Buffer reflection spectrum under pump. We represent the amplitude of the field reflected off the buffer as in Fig. 3c (top, curves offset for readability), together with the phase of the reflected signal (middle) and a representation in the Fresnel plane (bottom). The spectra are well reproduced by simulations based on the effective Hamiltonian (S35) (lines).
S4 Numerical simulations and fitting procedures
S4.1 Frequency maps
In this section, we detail how the circuit parameters are estimated by analyzing the dependency of each mode resonance frequency against DC bias fluxes, in the neighborhood of the ATS operating points (see Fig. 2a, operating points appearing as saddle points in the color map). Our model Hamiltonian is derived from the three-mode effective circuit pictured in Fig. 2a (obtained after Foster synthesis of the environment seen by the SQUID) and includes the non-idealities listed in Sec. S2, it reads
$\begin{split}{\bf H}=\sum_{i}\hbar\omega_{i}{\bf a}_{i}^{\dagger}{\bf a}_{i}&-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\bm{\varphi}-\varphi_{\Delta})-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\bm{\varphi}-\varphi_{\Delta})\\ &+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\bm{\varphi}-2\varphi_{\Delta})+\frac{E_{L}}{N^{2}}\big{(}\cos(\bm{\varphi})+\frac{1}{2}\bm{\varphi}^{2}\big{)}\end{split}$ (S19)
In this expression, $i\in\{a,b,r\}$ labels each mode, $\bm{\varphi}=\sum_{i}\varphi_{i}^{ZPF}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ is the reduced phase across the SQUID, $\varphi_{i}^{ZPF}=(\pi Z_{i}/R_{q})^{1/2}$ represents the vacuum fluctuations of the phase of each mode across the SQUID. We have defined the reduced common and differential bias fluxes $(\varphi_{\Sigma},\varphi_{\Delta})$ in Eq. (S7). As detailed in Sec. S2, we have allowed for the ATS junctions to be imbalanced with respective energies $E_{J_{1}}$ and $E_{J_{2}}$ and defined $E_{\Sigma}=E_{J_{1}}+E_{J_{2}}$ and $E_{\Delta}=E_{J_{1}}-E_{J_{2}}$. The last term on the first line stems from this imbalance (see Sec. S2). The first term on the second line stems from series inductances in the arms of the SQUID. We have assumed that series inductances are identical on both arms (inductive energy $\epsilon_{L}$) and defined $e_{\Sigma}=(E^{2}_{J_{1}}+E^{2}_{J_{2}})/(4\epsilon_{L})$ (a last term in $(E^{2}_{J_{1}}-E^{2}_{J_{2}})/(4\epsilon_{L})$ is neglected). Finally, the last term on the second line stems from the non-linear contribution of the superinductors potential (each a chain of $N=25$ for a total inductance $E_{L}$ [^8] This term is only valid in the RWA since the modes phase fluctuations can have opposed signs accross both inductors. However, the only contributions that survives the RWA do not depend on this sign.
In practice, we first record the frequency of the readout mode over a full period of modulation for both fluxes. After locating three saddle points on the recorded map, we find the linear transform relating DC currents applied to the circuit and reduced fluxes $\varphi_{\Sigma}$ and $\varphi_{\Delta}$. We then measure all three modes resonance frequencies nearby the saddle points. More specifically, the fluxes are varied along cuts intersecting $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,\pm\pi/2)$, and we record the transition frequencies $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ of the memory and the lowest transition frequency of the buffer and readout modes. In order to facilitate the fit of these datasets with the model Hamiltonian (S19), we first fit the measured frequencies along each cut with a parabola, and report the extracted curvature against cut angle in Fig. S12. We subsequently fit the curvature pattern with our model. We first detail a simplified classical analysis of the circuit around the frequency saddle points, which is the rationale for the fitting procedure to the full quantum model detailed in the following section.
Classical analysis of the frequency map curvature at the saddle points
We here compute simple analytical formula for the frequency of each mode against externally applied fluxes in the neighborhood of the saddle points, using a classical model. We neglect all circuit non-idealities ($E_{\Delta}=e_{\Sigma}=E_{L}/N^{2}=0$ in (S19)). Moreover, for simplicity, we consider two only modes $a$ and $b$ with phases $\varphi_{a}$ and $\varphi_{b}$. We will generalize the obtained formula to the three-mode case at the end of the calculation.
Letting $E_{L_{a}}$ and $E_{L_{b}}$ the modes inductive energies, and $E_{C_{a}}$ and $E_{C_{b}}$ their capacitive energies (related to the modes resonance frequencies and impedances by $\omega_{a,b}=(8E_{L_{a,b}}E_{C_{a,b}})^{1/2}$ and $Z_{a,b}=R_{q}(2E_{C_{a,b}}/E_{L_{a,b}})^{1/2}/\pi$), we express the circuit inductive potential in a coordinate system $(\epsilon_{\Sigma},\epsilon_{\Delta})=(\varphi_{\Sigma},\varphi_{\Delta})-(-\pi/2,-\pi/2)$ centered on the saddle point $\mathrm{SP}_{2}$ (see Fig. 2b):
$V(\varphi_{a},\varphi_{b})=\frac{1}{2}E_{L_{a}}\varphi_{a}^{2}+\frac{1}{2}E_{L_{b}}\varphi_{b}^{2}+E_{\Sigma}\sin(\epsilon_{\Sigma})\sin(\varphi_{a}+\varphi_{b}-\epsilon_{\Delta})$ (S20)
We consider cuts through the saddle point parametrized by an angle $\theta$
$\epsilon_{\Sigma}$ $=\epsilon\cos(\theta)$ (S21)
$\epsilon_{\Delta}$ $=\epsilon\sin(\theta)$
with $\epsilon$ varying on a range much smaller than 1. We then expand the potential to second order in $\epsilon$. Noting $c=\cos(\theta)$ and $s=\sin(\theta)$ for compactness, we obtain
$V(\varphi_{a},\varphi_{b})=$ $(\displaystyle\frac{1}{2}E_{L_{a}}\varphi_{a}^{2}+\frac{1}{2}E_{L_{b}}\varphi_{b}^{2}+E_{\Sigma}c\epsilon\Big{(}$S22)
$\qquad-\epsilon s\big{(}\cos(\varphi_{a})\cos(\varphi_{b})-\sin(\varphi_{a})\sin(\varphi_{b})\big{)}$
$\qquad+\sin(\varphi_{a})\cos(\varphi_{b})+\cos(\varphi_{a})\sin(\varphi_{b})$
[…]
$+\mathcal{O}(\epsilon^{3})$
The first step is to look for the coordinates $(\varphi_{a}^{m},\varphi_{b}^{m})$ that minimize the inductive energy. We will then expand the potential to second order to estimate the frequency of a classical particle oscillating around this minimum. The coordinates of the minimum are approximated to second order in $\epsilon$. Since $(\varphi_{a}^{m},\varphi_{b}^{m})=(0,0)$ when $\epsilon=0$, they are parametrized as
$\varphi_{a}^{m}$ $=\alpha_{a}\epsilon+\beta_{a}\epsilon^{2}+\mathcal{O}(\epsilon^{3})$ (S23)
$\varphi_{b}^{m}$ $=\alpha_{b}\epsilon+\beta_{b}\epsilon^{2}+\mathcal{O}(\epsilon^{3})$
Using $\frac{\partial V}{\partial\varphi_{a}}|_{(\varphi_{a}^{m},\varphi_{b}^{m})}=0$, we get
$\alpha_{a}\epsilon+\beta_{a}\epsilon^{2}+\epsilon r_{a}c+\mathcal{O}(\epsilon^{3})=0$ (S24)
where we have let $r_{a}=\frac{E_{\Sigma}}{E_{L_{a}}}$.
Figure S11: Effective circuit corresponding to the inductive potential encoded by the Hessian matrix (S27) and the charging energies found by Foster synthesis of the SQUID environment.
We can proceed similarly for the second variable. After identifying terms with same order in $\epsilon$ in the equations, we get
$\varphi_{a}^{m}$ $=-\epsilon r_{a}c+\mathcal{O}(\epsilon^{3})$ (S25)
$\varphi_{b}^{m}$ $=-\epsilon r_{b}c+\mathcal{O}(\epsilon^{3})$
We can inject these expressions in the formulas giving the second derivatives of the potential around the minimum
$\frac{\partial^{2}V}{\partial\varphi_{a}^{2}}|_{(\varphi_{a}^{m},\varphi_{b}^{m})}$ $=E_{L_{a}}\big{(}1+\epsilon^{2}r_{a}c(s+r_{a}c+r_{b}c)\big{)}+\mathcal{O}(\epsilon^{3})$ (S26)
$\frac{\partial^{2}V}{\partial\varphi_{b}^{2}}|_{(\varphi_{a}^{m},\varphi_{b}^{m})}$ $=E_{L_{b}}\big{(}1+\epsilon^{2}r_{b}c(s+r_{a}c+r_{b}c)\big{)}+\mathcal{O}(\epsilon^{3})$
$\frac{\partial^{2}V}{\partial\varphi_{a}\partial\varphi_{b}}|_{(\varphi_{a}^{m},\varphi_{b}^{m})}$ $=E_{\Sigma}\epsilon^{2}c(s+r_{a}c+r_{b}c)+\mathcal{O}(\epsilon^{3})$
which, letting $F=\epsilon^{2}E_{\Sigma}c(r_{a}c+r_{b}c+s)$, is compactly encoded by the Hessian matrix
$\begin{pmatrix}E_{L_{a}}+F&F\\ F&E_{L_{b}}+F\end{pmatrix}$ (S27)
at second order in $\epsilon$. Since the modes charging energies do not vary as a function of flux, this effective potential combined with the charging energies correspond to the circuit pictured in Fig. S11, where we have defined $L_{F}=\frac{\varphi_{0}^{2}}{F}$. We can write down the circuit Lagrangian and find the two dressed mode frequencies
$\omega_{\pm}^{2}=\frac{\tilde{\omega}_{a}^{2}+\tilde{\omega}_{b}^{2}}{2}\pm\frac{\big{(}(\tilde{\omega}_{a}^{2}-\tilde{\omega}_{b}^{2})^{2}+4\omega_{F}^{4}\big{)}^{1/2}}{2}$ (S28)
where we use the notations $\tilde{\omega}_{a}=\frac{1}{\sqrt{(L_{a}\parallel L_{F})C_{a}}}$, $\tilde{\omega}_{b}=\frac{1}{\sqrt{(L_{b}\parallel L_{F})C_{b}}}$, ${\omega}^{4}_{F}=\frac{1}{{L_{F}^{2}C_{a}C_{b}}}$.
At second order in $\epsilon$, $\omega_{F}^{4}=0$ and we find the mode frequencies to be
$\tilde{\omega}_{a}$ $=\omega_{a}\big{(}1+\frac{\epsilon^{2}}{2}r_{a}c(r_{a}c+r_{b}c+s)\big{)}$ (S29)
$\tilde{\omega}_{b}$ $=\omega_{b}\big{(}1+\frac{\epsilon^{2}}{2}r_{b}c(r_{a}c+r_{b}c+s)\big{)}$
Finally, we generalize this formula to our three-mode circuit to get each mode curvature $C_{i}$ ($i=a,b,r$) against the cut angle $\theta$ at the saddle-point
$C_{i}(\theta)=\omega_{i}r_{i}\cos(\theta)\big{(}(r_{a}+r_{b}+r_{c})\cos(\theta)+\sin(\theta)\big{)}$ (S30)
where we recall that $r_{i}=\frac{E_{\Sigma}}{E_{L_{i}}}$ and $\omega_{i}$ is the mode resonance frequency at the saddle point.
A few remarks are here in order.
  • One can obtain a formula for the curvature at the other saddle point located at $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,+\pi/2)$ (corresponding to $\mathrm{SP}_{1}$ in Fig. 2, up to a $2\pi$ offset in $\varphi_{\Delta}$) by substituting $r_{i}\rightarrow-r_{i}$
  • Once the modes resonance frequencies at the saddle points are known (directly measurable), the curvatures of the modes frequencies against cut angle fully characterize the frequency maps in the neighborhood of the saddle points. Therefore, measuring and fitting these curvatures is an efficient and straightforward way to extract circuit parameters. The method is not sufficient to unambiguously estimate all circuit parameters though, which requires one extra constraint (below, this constraint is given by the curvature of the $|0\rangle_{a}\rightarrow|3\rangle_{a}$ transition).
  • Imbalance of the SQUID junctions modifies the inductive Hessian matrix (S27) following
    $H=\begin{pmatrix}E_{L_{a}}-E_{\Delta}+F&-E_{\Delta}+F\\ -E_{\Delta}+F&E_{L_{b}}-E_{\Delta}+F\end{pmatrix}$ (S31)
    where $F$ can be expressed as a function of the ratios $E_{\Sigma}/E_{L_{i}}$ and $E_{\Delta}/E_{L_{i}}$, with a similar expression substituting $E_{\Sigma}\rightarrow-E_{\Sigma}$ and $E_{\Delta}\rightarrow-E_{\Delta}$ at the other saddle point. We find that the frequency splitting of each mode between the two saddle point reads $\delta\omega_{i}=\omega_{i}E_{\Delta}/E_{L_{i}}$ which allows us to estimate $E_{\Delta}$, of the order of $10^{-5}E_{\Sigma}$ in our experiment. This small imbalance is not expected to impact the curvatures at the saddle point derived above.
Fits to the full quantum model
Figure S12: Curvature of the frequency maps around the saddle points a-b) Measured (filled circles) and simulated (plain lines) curvature of the memory and buffer low-energy transition frequency at the two non-equivalent saddle points, against cut angle $\theta$ (defined in Eq. (S21)). Simulated values use the model Hamiltonian (S19) fitted to the data. Dashed lines represent the predictions from the classical model (S30) with the same parameters. These only reproduce qualitatively the data. Note that the curvature of the readout mode has been enhanced by a factor 20 for visibility. c) Predictions from the quantum and classical model for a hypothetical circuit whose modes have five times lower phase fluctuations. Here, the two models agree quantitatively.
In the classical model detailed above, we have neglected the vacuum fluctuations of each mode. Therefore, we do not expect the predictions to be accurate when the modes vacuum fluctuations approach 1, which is the case in our experiment. After a first coarse estimate of the circuit parameters using our classical model, we refine our analysis by numerically diagonalizing the Hamiltonian (S19). In Fig. S12a-b, we represent the curvature $C_{i}$ of each mode resonance frequency around the saddle point against cut angle $\theta$ (see Eq. (S21)), together with the result of these simulations (for the estimated circuit parameters) and the predictions from the classical analysis. The latter only qualitatively reproduce the data. For comparison, we represent in Fig. S12c the predictions from the Hamiltonian diagonalization against classical predictions for a circuit whose modes have five times smaller vacuum fluctuations across the SQUID. Here, the agreement is quantitative.
As stated above, measuring and fitting a single transition frequency for each mode is not sufficient to estimate independently the total Josephson energy of the SQUID $E_{\Sigma}$ and the value of the phase fluctuations $\varphi^{ZPF}_{i}$ of each across the SQUID (equivalently $E_{\Sigma}$ and each mode inductive energy $E_{L_{i}}$). Therefore, we additionally consider the dependency of the $|0\rangle_{a}\rightarrow|3\rangle_{a}$ transition frequency against flux biases, which provides an additional constraint on the estimated parameters. Intuitively, the curvature of the anharmonicity $\omega_{03}-3\omega_{01}$ at the saddle point is set by $E_{\Sigma}{\varphi_{a}^{ZPF}}^{4}$ (at leading order of the non-linearity). On the other hand, the curvature of $\omega_{01}$ depends on $E_{\Sigma}{\varphi_{a}^{ZPF}}^{2}$ only, such that both $E_{\Sigma}$ and ${\varphi_{a}^{ZPF}}$ can be fitted independently.
We next extract the energy difference between the two SQUID junctions $E_{\Delta}$ through the frequency splitting of the transition $|0\rangle_{a}\rightarrow|1\rangle_{a}$ at the two saddle points. As for the parameter $e_{\Sigma}$—accounting for the non-linearity induced by series inductance in the arms of the SQUID—it is estimated via the the value of $\omega_{03}-3\omega_{01}$ averaged over both saddle points. Given the small value of these two parameters, they only marginally impact the fits of the curvatures considered above. Finally we note that the Kerr non-linearity inherited from the circuit superinductances is also included in our numerical model but has no significant impact on the numerical results ($E_{L}$ is estimated from the nominal value of the superinductors).
The fitted values for all parameters are summarized in Table. 1.
S4.2 Dynamics under the four-to one photon exchange interaction
In this section, we detail our model for the memory and buffer dynamics under the pump microwave addressing the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$. We then detail the fitting procedure for the datasets presented in Fig. 3b-c.
Model dynamics
We start from the circuit Hamiltonian (S19), which is time-dependent for flux biases $(\varphi_{\Sigma},\varphi_{\Delta})=(-\frac{\pi}{2}+\epsilon(t),-\frac{\pi}{2})$ that include a small AC component ($\epsilon(t)\ll 1$). Neglecting the readout mode, expanding at first order in $\epsilon$ and including a direct linear drive on the buffer (needed to reproduce the buffer reflection spectra in Fig. 3c), we get
$\begin{split}{\bf H}(t)=&\hbar\omega_{a}{\bf a^{\dagger}}{\bf a}+\hbar\omega_{b}{\bf b^{\dagger}}{\bf b}+\epsilon(t)E_{\Sigma}\sin(\bm{\varphi})\\ &+E_{\Delta}\cos(\bm{\varphi})+e_{\Sigma}\cos(2\bm{\varphi})+\frac{E_{L}}{N^{2}}\big{(}\cos(\bm{\varphi})+\frac{1}{2}\bm{\varphi}^{2}\big{)}\\ &+\epsilon_{b}(t)({\bf b}+{\bf b^{\dagger}})\end{split}$ (S32)
where $\bm{\varphi}=\varphi_{a}^{ZPF}({\bf a}+{\bf a^{\dagger}})+\varphi_{b}^{ZPF}({\bf b}+{\bf b^{\dagger}})$. Our goal is to derive an effective static Hamiltonian allowing for numerical simulation. To this end, we move to a rotating frame for both modes, and select resonant or near-resonant terms.
We assume the pump to be near-resonant with the four-to-one photon transition $\epsilon(t)=\xi e^{i(4\tilde{\omega}_{a}-\tilde{\omega}_{b}+\Delta)t}+\mathrm{c.c}.$ and the linear drive to be near resonant with the buffer $\epsilon_{b}(t)=\xi_{b}e^{i(\tilde{\omega}_{b}+\delta)t}+\mathrm{c.c}.$ with $\Delta,\delta\ll\tilde{\omega}_{a},\tilde{\omega}_{b}$. Here, $\tilde{\omega}_{a,b}$ denote the modes resonance frequencies dressed by the pump, which are assumed to be perfectly calibrated by the four-step procedure detailed in Sec. S3.2. When simulating the time-decay of the state $|4\rangle_{a}$, we thus set $\Delta=0$ and $\xi_{b}=0$ (pump on resonance, no buffer drive, see Fig. 3b). When reproducing the buffer reflection spectra (Fig. 3c), $\delta$ is the detuning between the applied buffer drive and the buffer resonance frequency dressed by the pump. We give a non-zero value to $\Delta=\omega_{b_{3}}-\omega_{b_{1}}$ to account for the slightly off-resonant pump (we use the convention of the four-step calibration procedure and recall that the spectra are recorded at step 3). In practice, this small detuning has no visible impact on the simulation results.
We now move to frames rotating with both the pump and the linear drive
$\begin{split}a&\rightarrow a^{-i\big{(}\tilde{\omega}_{a}+\frac{\Delta+\delta}{4}\big{)}t}\\ b&\rightarrow b^{-i(\tilde{\omega}_{b}+\delta)t}.\end{split}$ (S33)
First considering the terms on the second line of the Hamiltonian (S32) stemming from non-idealities of the circuit, we discard contributions that are fast-oscillating in the rotating frame, i.e. which do not preserve photon number. This amounts to keeping only diagonal terms in the Fock basis and we end up with a diagonal Hamiltonian
${\bf H}_{\chi}=\sum_{n,m}\chi_{nm}|n\rangle_{a}|m\rangle_{b}\langle n|_{a}\langle m|_{b}$ (S34)
where the dispersive shifts $\chi_{nm}$ are computed from the non-driven version of Hamiltonian (S32) (setting $\epsilon=\epsilon_{b}=0$). In detail, we first diagonalize the non-driven Hamiltonian and label $(n,m)$ each eigenstate by identifying it to the Fock state $|n\rangle_{a}|m\rangle_{b}$ it maximally overlaps with  [^9] The maximum overlap (defined for two states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ as $|\langle\psi_{1}|\psi_{2}\rangle|^{2}$) is larger than 0.9995 for all $(n,m)$. Next, we label each state energy $E_{n,m}$ after offsetting the energy scale such that $E_{0,0}=0$. Finally, we find the dispersive shift $\chi_{nm}=E_{n,m}-n\frac{E_{4,0}}{4}-mE_{0,1}$. We note that with our definitions, the dispersive Hamiltonian ${\bf H}_{\chi}$ does not contribute to the energy splitting between $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$, which is expected given our choice of rotating frames. We also note that by computing ${\bf H}_{\chi}$ from the non-driven Hamiltonian (S32), we have neglected a small renormalization by the pump. In practice, the exact value of ${\bf H}_{\chi}$ does not significantly impact numerical simulations in the regime we are in, as detailed below.
Now turning to the term $\epsilon(t)E_{\Sigma}\sin(\bm{\varphi})$, we get in the RWA a resonant four-to-one photon exchange interaction at rate $g_{4}=\frac{E_{\Sigma}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ (see Sec. S1.1). Overall, we get the effective static Hamiltonian
${\bf H}_{RWA}=-\frac{\Delta+\delta}{4}{\bf a}^{\dagger}{\bf a}-\delta{\bf b}^{\dagger}{\bf b}+{\bf H}_{\chi}+g_{4}({\bf a}^{\dagger 4}{\bf b}+{\bf a}^{4}{\bf b^{\dagger}})+\xi_{b}{\bf b}+\xi_{b}^{\ast}{\bf b^{\dagger}}$ (S35)
Finally, we model dissipation and dephasing in our system by considering the following Lindblad dissipators
  • $\gamma_{k}\mathcal{D}[|k-1\rangle\langle k|]$ for $1\leq i\leq 4$. These dissipators account for the relaxation of population in state $|k\rangle$ through single-photon emission. The rates $\gamma_{k}$ for $k\leq 3$ are independently calibrated for each pump amplitude (see Sec. S3.3). Since we do not consider any other decay mechanism from these states, they are identical to the rates $\Gamma_{k}$ reported in Fig. 4. Note that their values do not follow the expected linear scaling in $k$, preventing us from modeling single-photon dissipation with a single dissipator $\mathcal{D}[{\bf a}]$. Given the memory anharmonicity, this behavior is expected if the coupling to the bath in which single-photons are emitted—or the bath density of states—is not constant in frequency. As for $\gamma_{4}$, it cannot be unambiguously estimated for a non-zero pump amplitude since decay from state $|4\rangle_{a}$ under pump is dominated by the engineered four-photon dissipation channel. We therefore assume $\gamma_{4}$ to be constant over all considered pump amplitudes.
  • $\kappa_{b}\mathcal{D}[{\bf b}]$. This dissipator accounts for single-photon dissipation in the buffer. Here, we assume $\kappa_{b}$ to be constant for all pump amplitudes as radiative decay through the charge line is the dominant contribution to single-photon relaxation. Moreover the buffer is negligibly excited beyond the first excited state in our experiment, such that we do not need to adjust the relaxation rate for higher excited states in our model. $\kappa_{b}$ is independently calibrated by fitting the buffer reflection spectra when the pump is off.
  • $2\kappa_{\phi}\mathcal{D}[{\bf a^{\dagger}}{\bf a}]$. This dissipator accounts for pure dephasing of the memory. The dephasing rate is fitted together with the datasets presented in Fig. 3 and is in good agreement with an independent measurement of the dephasing of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition. Note that this rate is enhanced by the application of the off-resonant pump employed to shift the memory transitions away from spurious TLS resonances (see Sec. S5) and was seen to fluctuate over the course of hours.
Fitting procedure
Our goal here is to fit the datasets presented in Fig. 3b-c with the model presented above. The fit parameters are
  • The value of $g_{4}$ for each pump amplitude. The fitted values are reported in 3d.
  • The displacement rate $\xi_{b}$ entailed by the linear drive on the buffer. The fitted value is $\xi_{b}=2\pi\times 175~{}$kHz.
  • The memory pure dephasing rate $\kappa_{\phi}$ (value reported in Table. 1).
  • A proportionality factor between the differential readout signal $S_{4}$ and the occupation of state $|4\rangle_{a}$ (determining the y-scale in Fig. 3b).
  • A scaling factor accounting for the cryostat input and output lines attenuation (determining the amplitude of off-resonant waves reflected from the buffer in Fig. 3c and Fig. S10).
In order to limit computational costs when fitting, we simplify our model based on the exact experiment that needs to be reproduced numerically. Before detailing and justifying these simplifications, we note that they are taken solely when fitting experimental data to extract unknown parameter values. Using the fitted values, we then run the model described above with no simplifying assumption to produce the simulated curves presented in Fig. 3.
Figure S13: Simplified six-level model employed to fit the datasets presented in Fig. 4. The pump (plain purple arrow) and the linear drive applied to the buffer (plain orange arrow) are each resonant with a single transition only (detuned transitions represented by dashed arrows). As a result, the system state is confined to a six-level manifold (levels pictured in black) and never visits other states (levels pictured in gray).
We first notice that for both datasets presented Fig. 3b-c, the system state is constrained to a 6-dimension manifold spanned by $\{|0\rangle_{a}|0\rangle_{b},|1\rangle_{a}|0\rangle_{b},|2\rangle_{a}|0\rangle_{b},|3\rangle_{a}|0\rangle_{b},|4\rangle_{a}|0\rangle_{b},|0\rangle_{a}|1\rangle_{b}\}$ (see Fig. S13). Indeed, the pump (purple arrow) is only resonant with the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition and is weak enough that no other transition can be driven ($g_{4}\sqrt{\frac{(n+4)!}{n!}(m+1)}\ll|(\chi_{n+4,m}-\chi_{n,m+1})|$ for $n>0$ or $m>0$). Similarly, the linear drive on the buffer (only on when recording the reflection spectra in Fig. 3c) is resonant with the $|0\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition only ($\xi_{b}\sqrt{m+1}\ll|(\chi_{n,m+1}-\chi_{n,m})|$ for $n>0$ or $m>0$).
For the data presented in Fig. 3b, we can further simplify our model by noticing that when the system relaxes from state $|4\rangle_{a}|0\rangle_{b}$ (either though single-photon relaxation to $|3\rangle_{a}|0\rangle_{b}$ at rate $\gamma_{4}$ or through a four-to-one photon exchange driven by the pump), it never returns to $|4\rangle_{a}|0\rangle_{b}$. Since we are only interested in computing the decay rate from $|4\rangle_{a}|0\rangle_{b}$, we can consider an effective three-level system $\{|D\rangle,|A\rangle,|B\rangle\}$ with $|A\rangle=|4\rangle_{a}|0\rangle_{b}$, $|B\rangle=|0\rangle_{a}|1\rangle_{b}$ and $|D\rangle$ a dark state to which $|A\rangle$ relaxes at rate $\gamma_{4}$ and $|B\rangle$ relaxes at rate $\kappa_{b}$. Finally, we use the Lindblad-adjoint formalism [27] to encode the evolution of the expectation value of the operators $|A\rangle\langle A|$, $|B\rangle\langle B|$, $|A\rangle\langle B|$, and $|B\rangle\langle A|$ in a $4\times 4$ evolution matrix. Note that beside allowing fast simulations of the evolution of the expectation value of $|A\rangle\langle A|$ (used to fit the data in Fig. 3b), we also perform spectral analysis of this matrix to extract the relaxation rate $\Gamma_{4}$ corresponding to the matrix eigenvalue closest to zero (reported in Fig. 4).
For the data presented in Fig. 3c, reduction of the system Hilbert space is less obvious. Here, we are interested in the steady-state value of $\langle{\bf b}\rangle$. We directly apply the Lindblad adjoint formalism to encode the evolution $\langle{\bf b}\rangle$ along with the expectation value of ten other orthogonal operators (which we compute numerically and do not detail here) in a $11\times 1$ 11 matrix $M$. The steady-state is directly found from the single vector in the kernel of $M$.
With these model simplifications, we perform a global fit of the two datasets in Fig. 3b-c with a simple least-square method. As a sanity check of our simplifying assumptions, we finally perform a full Lindblad master equation featuring the effective Hamiltonian ${\bf H}_{RWA}$ and the dissipators $\kappa_{b}\mathcal{D}[{\bf b}]$, $2\kappa_{\phi}\mathcal{D}[{\bf a^{\dagger}}{\bf a}]$ and $\gamma_{k}\mathcal{D}[|k-1\rangle\langle k|]$ for $1\leq i\leq 4$, with parameter values extracted from the above fit. The results of these simulations are plotted in Fig. 3b-c and Fig. S10 and quantitatively reproduce our data.
S5 Tuning the system away from TLS resonances
In our experiment, some of the memory transitions—either single-photon or three-photon transitions—were transiently observed to shift or split. Typically, a given transition that would yield a clean Lorentzian response in two-tone spectroscopy would suddenly yield a two-peak response. This behavior was either be permanent (until the system was cycled around 20 K before being cooled back down at base temperature) or appearing on and off on a minute timescale (see Fig. S14). The buffer lowest energy transition was also witnessed to split at one occasion.
We attribute this effect to hybridization of the system to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [57, 58]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, the probability that at least one of the memory or buffer transitions be affected by the TLS bath was high enough that waiting for the bath to reconfigure in a more favorable configuration (either passively or cycling the system to room temperature) was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the circuit transitions away from any resonant coupling with TLS’s.
Various mechanisms can contribute to a transition frequency shift proportional to the power of an off-resonant drive. In particular, a flux drive applied to the ATS loops with both a symmetric and an anti-symmetric component yields such a shift (see Sec. S2). The Stark effect [78] is another possible mechanism owing to the modes non-linearity and given that their states are linearly displaced by the drive. Such a displacement is expected both for a symmetric flux drive (through the leading term in the ATS sine Hamiltonian) and for an anti-symmetric flux drive (through induction in the circuit loops [75]). This latter mechanism is also expected to degrade the system coherence through measurement-induced dephasing [78]. Measurement-induced transitions are also a concern for some specific drive frequencies [59, 60, 54, 61]. While we did not quantitatively investigate these mechanisms in our experiment, the fact that a frequency shift of the circuit transitions is expected at any flux drive frequency lets us select the one which minimizes decoherence.
In detail, we use the characteristic decay time of Rabi oscillations driven on the memory $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition as a proxy for the system coherence. As illustrated in Fig. S15, we sweep the frequency of an off-resonant drive over a few GHz range. For each frequency, we adjust the drive power to get a 17 MHz shift of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition (which is approximately the value used in our experiment) and record Rabi oscillations in these conditions. We then select the off-resonant drive frequency yielding the longest-lived Rabi oscillations (at 9.25 GHz).
When acquiring the datasets presented in Fig. 3b-c, we apply this off-resonant drive at a power of 5.3 dBm, referenced as in Fig. S15 (at the output of the microwave source labeled Stark shifting pump in Fig. S3). The drive is turned on during the system preparation, application of the four-photon dissipation channel and the measurement. In these conditions, the $T_{1}$ of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition is measured to be of 7.5 $\mu$s (see Fig. S9 for a null pump amplitude) and in a later measurement, we observed the Ramsey free-induction decay time to be of 3.5 $\mu$s, corresponding to a pure dephasing rate of $2\pi\times 35$ kHz. This is in good agreement with the value fitted from the datasets presented in Fig. 3b-c. These lifetimes are slightly degraded when compared to those measured in absence of the off-resonant drive ($T_{1}\lesssim 11~{}\mu$s, $T_{2}\lesssim 6~{}\mu$s), even though these were seen to fluctuate over the course of days.
Figure S14: Signature of hybridization to a Two-Level System. Response of the system when probed in two-tone spectroscopy around the resonance frequency of the $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ transition (driven through the flux lines). The two traces are taken 8 minutes apart. On the later one (orange), the line is split in two peaks of approximately equal height.
Figure S15: Shifting the system transitions out of resonance from TLS’s. A controlled shift of all transitions is realized by the application of a strong off-resonant drive (labeled Stark shifting pump in Fig. S3). In order to select the drive frequency and power minimizing decoherence, we sweep the drive frequency and adjust its power to reach a Stark shift of 17 MHz for the memory $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition (corresponding power referenced at the microwave source output port reported in the bottom panel). We then record Rabi oscillations while this off-resonant drive is on (top panel), and select the drive frequency for which the oscillations are longest-lived (at 9.25 GHz).